### v1

# FEDONet : Fourier-Embedded DeepONet for Spectrally Accurate Operator Learning

Abstract
Deep Operator Networks (DeepONets) have recently emerged as powerful data-driven frameworks for learning nonlinear operators, particularly suited for approximating solutions to partial differential equations (PDEs). Despite their promising capabilities, the standard implementation of DeepONets, which typically employs fully connected linear layers in the trunk network, can encounter limitations in capturing complex spatial structures inherent to various PDEs. To address this, we introduce Fourier-embedded trunk networks within the DeepONet architecture, leveraging random Fourier feature mappings to enrich spatial representation capabilities. Our proposed Fourier-embedded DeepONet, FEDONet demonstrates superior performance compared to the traditional DeepONet across a comprehensive suite of PDE-driven datasets, including the two-dimensional Poisson equation, Burgers’ equation, the Lorenz-63 chaotic system, Eikonal equation, Allen–Cahn equation, Kuramoto–Sivashinsky equation, and the Lorenz-96 system. Empirical evaluations of FEDONet consistently show significant improvements in solution reconstruction accuracy, with average relative $\ell_{2}$ performance gains ranging between 2-3x compared to the DeepONet baseline. This study highlights the effectiveness of Fourier embeddings in enhancing neural operator learning, offering a robust and broadly applicable methodology for PDE surrogate modeling.
1 Introduction
Partial differential equations (PDEs) underpin models ranging from turbulent-flow dynamics and heat diffusion to electromagnetics and biological transport. Classical numerical solvers—finite-difference, finite-element and spectral methods—provide high-fidelity solutions but incur a prohibitive computational burden when faced with high-dimensional, multiscale or stiff systems [1, 2, 3]. The quest for rapid surrogates has a long pedigree: universal-approximation theorems for neural networks [4, 5, 6, 7], reduced-order techniques based on proper-orthogonal decomposition (POD) [8, 9], kernel–Gaussian-process surrogates [10], and mesh-free radial-basis schemes [11, 12]. Yet, translating these ideas into *function-to-function* (operator) mappings that respect arbitrary discretisations remained elusive.
The modern wave of operator learning addresses this gap by training neural networks that act directly on infinite-dimensional function spaces and generalise across grid resolutions. Deep Operator Networks (DeepONets) [13] and the broader Neural Operator family [14, 15, 16] establish a branch–trunk architecture with rigorous operator–universal-approximation guarantees. Fourier Neural Operators (FNOs) extend this paradigm through global spectral convolutions, delivering impressive speed-ups on parametric PDEs [17]. Subsequent variants exploit multiwavelets [18, 19], convolutional kernels [20], multi-grid tensorisation [21, 22] and geometry-aware deformations [23] to boost resolution or handle complex domains.
Beyond kernel choices, researchers have tackled discontinuities [24], derivative supervision [25], irregular meshes via graph trunks [26], active multi-resolution sampling [27], and U-Net/wavelet hybrids [28, 29]. Random Fourier features have proved especially effective at encoding high-frequency structure in standard deep nets [30], yet remain under-explored inside operator networks whose trunk is typically a plain fully connected multilayer perceptron (MLP).
To incorporate governing laws and uncertainty, Physics-Informed Neural Networks (PINNs) [31, 32] evolved into Physics-Informed Neural Operators (PINO) [33, 34] and their variational or pseudo-physics relatives [35, 36]. Complementary efforts embed operators in latent or invertible spaces for data efficiency and bidirectional inference [37, 38, 39, 40].
Transformer-style attention has recently been adopted for operator learning [41, 42, 43, 44, 45] and for geometry generalisation [46]. Such architectures underpin landmark climate-scale surrogates such as FourCastNet [47] and its amortised FNO successor [48].
Despite the rich taxonomy above, the trunk of conventional DeepONets, charged with encoding spatial/temporal coordinates, remains a shallow MLP. This design struggles to capture the oscillatory or multi-scale patterns ubiquitous in nonlinear PDE solutions, leading to sub-optimal sample efficiency and accuracy. Recent trunk alternatives (e.g. graph [26] or wavelet U-Net [28]) introduce architectural complexity or domain-specific preprocessing.
We propose a Fourier-Embedded DeepONet (FEDONet) that replaces the raw coordinates fed into the trunk with a compact random Fourier-feature map [30]. The embedding supplies an implicit spectrum of sinusoidal basis functions, endowing the trunk with a richer hypothesis space at negligible parameter cost. FEDONet preserves the original branch–trunk separation, is plug-and-play with existing DeepONet codebases, and adds no runtime overhead beyond a single matrix multiplication.
  1. (i)
    We formulate the Fourier embedding of trunk coordinates, derive its effect on the operator neural tangent kernel, and show that it strictly enlarges the approximation class of DeepONets.
  2. (ii)
    We deliver a thorough benchmark across seven PDE families—steady Poisson, viscous Burgers, Allen–Cahn, Kuramoto–Sivashinsky, Eikonal, Lorenz–63 and Lorenz–96—demonstrating 48.93% average reduction in relative $\ell^{2}$ error versus a tuned vanilla DeepONet.
  3. (iii)
    We analyse hyperparameter sensitivity (mapping size and scale) and empirically confirm that the Fourier trunk improves convergence speed without sacrificing discretisation invariance.
By injecting a principled spectral bias into the DeepONet trunk, FEDONet closes the performance gap to frequency-domain operators such as FNO while retaining the flexibility and locality advantages of the original architecture.
2 Methodology
We aim to develop a data-driven framework that approximates nonlinear operators mapping between infinite-dimensional function spaces. Let $\Omega\subset\mathbb{R}^{D}$ be a bounded domain, and define the input and output function spaces:
$\mathcal{U}$ $=\left\{u:\mathcal{X}\rightarrow\mathbb{R}^{d_{u}}\right\},\quad\mathcal{X}\subseteq\mathbb{R}^{d_{x}},$ (1)
$\mathcal{S}$ $=\left\{s:\mathcal{Y}\rightarrow\mathbb{R}^{d_{s}}\right\},\quad\mathcal{Y}\subseteq\mathbb{R}^{d_{y}},$ (2)
where $\mathcal{U}$ denotes input functions (e.g., boundary or initial conditions) and $\mathcal{S}$ represents output fields (e.g., PDE solutions). Our goal is to learn an approximation $\mathcal{G}_{\theta}:\mathcal{U}\rightarrow\mathcal{S}$ to an unknown operator $\mathcal{G}$, where $\theta\in\Theta$ are trainable parameters.
In practice, we are given a dataset $\mathcal{D}=\{(u^{i},s^{i})\}_{i=1}^{N}$, comprising function pairs sampled from $\mathcal{G}$. DeepONet is a neural operator architecture well-suited to this task. It decomposes the learning problem into two components: Branch network $B_{\theta}$ that encodes discrete samples of $u$ at sensor points $\{x_{1},\dots,x_{m}\}$, Trunk network $T_{\theta}$ that takes in spatial (or spatiotemporal) coordinates $\zeta=(x,y,z,t)$.
The predicted value at location $\zeta$ is computed via an inner product:
$\mathcal{G}_{\theta}(u)(\zeta)=B_{\theta}(u)\cdot T_{\theta}(\zeta).$ (3)
2.1 Fourier Embeddings as Spectral Preconditioners
Standard MLPs exhibit a well-known spectral bias — they learn low-frequency components faster than high-frequency ones [49]. This bottleneck is particularly harmful in operator learning problems involving sharp gradients, multiscale dynamics, or oscillatory fields.
To address this, we propose injecting fixed Fourier embeddings into the trunk network input. Given a coordinate $\zeta\in\mathbb{R}^{d}$, we define a randomized Fourier feature map:
$\phi(\zeta)=\left[\sin(2\pi B\zeta),\cos(2\pi B\zeta)\right],\quad B_{ij}\sim\mathcal{N}(0,\sigma^{2}),$ (4)
where $B\in\mathbb{R}^{M\times d}$ is a Gaussian frequency matrix. Instead of feeding raw coordinates to the trunk, we input $\phi(\zeta)$ — enriching its spectral content.
The DeepONet prediction then becomes:
$\mathcal{G}_{\theta}(u)(\zeta)=B_{\theta}(u)\cdot T_{\theta}(\phi(\zeta))=\sum_{k=1}^{p}b_{k}(u)\cdot t_{k}(\phi(\zeta)).$ (5)
Figure 1: FEDONet: Fourier Embedded Deep operator Network
This transformation can be interpreted as a randomized kernel approximation:
$k(\zeta,\zeta^{\prime})\approx\phi(\zeta)^{\top}\phi(\zeta^{\prime}),$ (6)
effectively lifting $\zeta$ into a high-dimensional basis of sinusoidal functions. By Bochner’s theorem [50], such embeddings approximate shift-invariant kernels — implicitly kernelizing the trunk network. This expands its representational capacity, enabling better learning of high-frequency targets.
Beyond expressivity, Fourier embeddings improve the conditioning of the learning process. Under mild assumptions, the embedded features exhibit approximate whitening ( see A ) :
$\mathbb{E}_{\zeta}[\phi(\zeta)\phi(\zeta)^{\top}]\approx I,$ (7)
reducing feature correlations and leading to a more isotropic Neural Tangent Kernel (NTK) spectrum. This in turn improves optimization dynamics, stabilizes gradient flow, and accelerates convergence during training.
Let $\mathcal{H}_{\phi}$ denote the function class learnable with Fourier features, and $\mathcal{H}_{\text{vanilla}}$ the class learnable with raw coordinates. Then:
$\mathcal{H}_{\phi}\supset\mathcal{H}_{\text{vanilla}},$ (8)
with the inclusion being strict in problems with rich frequency content (see B). Fourier embeddings thus act as spectral preconditioners that align the input geometry with the task’s spectral demands.
2.2 Training Objective
We adopt a supervised learning framework to train the proposed FEDONet model. The objective is to approximate a target operator $G$ by minimizing the prediction error over a dataset of paired input–output functions. Each dataset element consists of a function $u^{(i)}\in\mathcal{U}$ and corresponding evaluations of the output $s^{(i)}_{j}=G(u^{(i)})(\zeta^{(i)}_{j})$ at query locations $\zeta^{(i)}_{j}\in\mathcal{Y}$.
Let the dataset be denoted by
$\mathcal{D}=\left\{\left(u^{(i)},\left\{\zeta^{(i)}_{j},s^{(i)}_{j}\right\}_{j=1}^{Q}\right)\right\}_{i=1}^{N},$
where $N$ is the number of input functions and $Q$ is the number of query points per function.
The model prediction at location $\zeta$ is given by:
$G_{\theta}(u)(\zeta)=\sum_{k=1}^{p}b_{k}(u)\cdot t_{k}(\phi(\zeta)),$
where $b_{k}(u)$ and $t_{k}(\phi(\zeta))$ denote the outputs of the branch and Fourier-embedded trunk networks, respectively.
The training objective is to minimize the empirical risk using the mean squared error (MSE) loss:
$\mathcal{L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{Q}\sum_{j=1}^{Q}\left\|G_{\theta}(u^{(i)})(\zeta^{(i)}_{j})-s^{(i)}_{j}\right\|^{2}.$
Model parameters are optimized using mini-batch stochastic gradient descent with the Adam optimizer. Learning rate schedules are selected based on the problem characteristics to ensure stable and efficient convergence. Each mini-batch consists of randomly sampled function–query pairs $(u,\zeta)$, enabling scalable and efficient training across diverse domains.
Notably, the Fourier embedding layer $\phi(\zeta)$ is fixed and non-trainable throughout training. This design choice ensures low overhead while enhancing model expressivity through spectral enrichment. Nonetheless, the architecture can be extended to incorporate learnable embeddings $\phi_{\theta}(\zeta)$, enabling adaptive spectral tuning tailored to problem-specific frequency content.
2.3 Evaluation and Generalization
After training, the FEDONet model is evaluated on unseen input functions $u_{\text{test}}$. To assess model performance, we report both quantitative and qualitative metrics:
  • Relative $L^{2}$ Error: Defined as
    $\varepsilon_{L^{2}}=\frac{\|G_{\theta}(u)-G(u)\|_{2}}{\|G(u)\|_{2}},$
    where the norm is computed over a dense evaluation grid. This metric quantifies the normalized discrepancy between predicted and reference output fields.
  • Spectral Fidelity: We perform Fourier analysis of the predicted and true fields to evaluate preservation of high-frequency content. Specifically, we compare the angle-integrated power spectra $E(k)$ across wavenumbers $k$, assessing the model’s ability to capture multiscale dynamics and resolve fine-scale features.
The ability to evaluate the learned operator at arbitrary coordinates makes FEDONet particularly suitable for scientific computing tasks requiring continuous surrogate modeling. In all benchmark scenarios, FEDONet demonstrates strong generalization, enhanced spectral accuracy, and robust performance across a wide range of PDE systems with varying stiffness, nonlinearity, and complexity.
3 Results
3.1 2D Poisson Equation
We created a synthetic dataset comprising numerical solutions of the two-dimensional (2D) Poisson equation, a fundamental PDE commonly employed in modeling steady-state processes such as heat conduction, electrostatics, and diffusion phenomena. The mathematical formulation of the Poisson equation is given by:
$\nabla^{2}u(x,y)=f(x,y),\quad(x,y)\in[0,1]^{2},$ (9)
accompanied by homogeneous Dirichlet boundary conditions:
$u(x,y)\big{|}_{\partial\Omega}=0.$ (10)
In this context, $u(x,y)$ denotes the solution field, while $f(x,y)$ represents a specified source term or coefficient.
To generate the dataset, we synthesized $N=10,000$ distinct instances of the coefficient function $f$ over a uniform spatial grid of resolution $128\times 128$. Each instance of the coefficient field was produced using Gaussian Random Fields (GRFs) characterized by specific spectral parameters, notably $\alpha=3$ and $\tau=3$, resulting in:
$f\sim\text{GRF}(\alpha=3,\tau=3).$ (11)
Once these coefficient fields were created, we numerically solved the associated Poisson equation to obtain the corresponding solution fields $u(x,y)$. The numerical discretization employed a finite difference approach, utilizing a conventional five-point stencil approximation of the Laplacian operator. By discretizing the spatial domain $\Omega=[0,1]^{2}$ into a $128\times 128$ structured grid, the PDE was transformed into a linear algebraic system:
$A\mathbf{u}=b,$ (12)
where $\mathbf{u}\in\mathbb{R}^{16384}$ denotes the vectorized solution field and $b\in\mathbb{R}^{16384}$ is the vector representation of the source term $f$. The discretized Laplacian matrix $A$ is defined explicitly as:
[…]
with homogeneous Dirichlet boundary conditions enforced directly by setting boundary rows of $A$ as identity matrices and corresponding entries in $f$ to zero.
Ultimately, this methodology yielded 10,000 independent pairs of coefficient-source fields and their respective solutions, forming a robust dataset suitable for training, validating, and benchmarking neural operator models. The dataset facilitates a comprehensive evaluation of the trade-offs in computational efficiency, accuracy, and reconstruction performance of our proposed neural operator architectures.
Figure 2 presents a representative test sample from the two–dimensional Poisson dataset. The input forcing field is shown on the left, followed by the exact solution, the prediction of a vanilla DeepONet, and the corresponding point-wise error. The bottom row repeats the comparison for our Fourier-Embedded DeepONet (FEDONet).
  • Spatial reconstruction. FEDONet reproduces the lobed extrema, smooth gradients and corner behaviour of the ground truth almost indistinguishably, whereas the vanilla network exhibits visible blurring and slight spatial shifts of the high-magnitude regions. The residual of the vanilla model oscillates between $\pm 1.5\times 10^{-3}$, displaying high-frequency artefacts; FEDONet reduces this range to $\pm 3\times 10^{-4}$, a five-fold improvement that manifests as a low-amplitude, coherent pattern.
  • Energy spectra. Figure 3 plots the angle-integrated energy spectrum $E_{k}$ for four randomly chosen hold-out fields (samples 29, 15, 24 and 17). Both networks match the reference spectrum at large scales ($k\lesssim 5$), but qualitative differences arise in the inertial and dissipation ranges:
    1. (a)
      The vanilla DeepONet (purple) consistently over-estimates energy for $k\gtrsim 8$, producing a spectral “pile-up” that grows by nearly an order of magnitude in sample 24—mirroring the spurious small-scale oscillations seen in its error map.
    2. (b)
      FEDONet (red) tracks the ground truth (dark blue) almost perfectly up to the Nyquist limit, deviating by at most a factor of two in the far dissipation tail and thereby preserving the correct energy cascade.
Figure 2: Model reconstruction comparison for the 2D Poisson Dataset.
Together, the spatial and spectral diagnostics confirm that Fourier embeddings in the trunk endow DeepONets with the high-frequency resolution and stability required for accurate operator learning: FEDONet not only slashes point-wise errors but also eliminates the artificial energy plateau that plagues fully-connected trunks.
Figure 3: Energy spectrum comparison between the ground truth, vanilla and Fourier DeepOnet reconstructions for the 2D Poisson Dataset.
3.2 Burgers’ Equation
The Burgers’ equation is a well-known 1D viscous equation. It has been studied extensively for DeepONets in prior literature. The equation is given by:
$\frac{\partial u}{\partial t}(x,t)+u\frac{\partial u}{\partial x}(x,t)=\nu\frac{\partial^{2}u}{\partial x^{2}}(x,t),\>\>\>\>\>\forall(x,t)\in[0,1]\times[0,1]$ (14)
where $x$ and $t$ denote the spatio-temporal coordinates and $\nu$ is the kinematic viscosity. We set $\nu=0.01$ along with initial and periodic boundary conditions as:
$\begin{split}\displaystyle u(x,0)=&\>s(x),\>\>\>\>\>\>\>\>\>\>\>\forall\>x\in[0,1]\\ u(0,t)=&\>u(1,t),\>\>\>\>\>\>\>\>\forall\>t\in[0,1]\\ \frac{\partial u}{\partial x}(0,t)=&\>\frac{\partial u}{\partial x}(1,t)\>\>\>\>\>\>\forall\>t\in[0,1]\end{split}$ (15)
where the initial condition $s(x)$ is sampled from a Gaussian Random Field, $s(x)\sim\mathcal{N}\left(0,25^{2}(-\Delta+5^{2}I)^{-4}\right)$, satisfying periodicity. The objective is to learn the nonlinear operator $\mathcal{G}$ mapping $s(x)$ to the full spatio-temporal solution $u(x,t)$ using both vanilla DeepONet and the proposed FEDONet.
Figure 4 shows the relative $\ell_{2}$ errors over 128 unseen test samples. FEDONet consistently outperforms DeepONet, with a mean error reduction from 4.01% to 2.63%, and a narrower spread in error distribution indicating increased reliability.
Figure 4: Burgers Relative $\ell_{2}$ error across 128 unseen samples for DeepONet and FEDONet.
Figure 5 presents the best-performing test sample, showing the ground truth solution alongside predictions from both models. While DeepONet captures the coarse solution structure, it struggles with steep gradients and exhibits visible smoothing near shock regions. In contrast, FEDONet produces significantly sharper and more accurate reconstructions, faithfully capturing shock fronts and fine-scale dynamics.
Figure 5: Best-performing test sample: Ground truth vs. DeepONet vs. FEDONet prediction.
The energy spectra of the predicted and ground truth fields, shown in Figure 6, further emphasize this advantage. FEDONet accurately captures the spectral decay across a wide range of wavenumbers, aligning with the physical expectations of dissipative transport phenomena. DeepONet, however, underestimates energy at higher modes, reflecting its inability to resolve fine features and spectral richness. This demonstrates that the Fourier embeddings effectively bridge the model’s spectral bias, enabling it to better represent high-frequency dynamics essential to Burgers-type flows.
These results underscore the importance of aligning model inductive bias with the physical spectrum of the target PDE. The enhanced performance of FEDONet in both physical and spectral domains highlights its potential as a more robust and generalizable operator learner for convection-dominated systems.
Figure 6: Temporal Evolution of Pointwise Error and Energy Spectrum: Best-Performing Sample vs. Ground Truth
Figure 7 shows the convergence of training losses for different batch sizes. FEDONet consistently converges faster and to lower loss values across all training configurations, indicating improved optimization characteristics and stability during training.
Figure 7: Training loss convergence for the 1D Burgers’ Equation: Comparison of Vanilla and Fourier Embedded DeepONets trained with varying batch sizes.
3.3 Lorenz–63 System
The Lorenz–63 system is a widely studied chaotic dynamical system that serves as a benchmark for evaluating the long-term stability and predictive accuracy of surrogate models. It is governed by the following system of coupled nonlinear ordinary differential equations:
$\frac{dx}{dt}$ $=\sigma(y-x),$ (16)
$\frac{dy}{dt}$ $=x(\rho-z)-y,$
$\frac{dz}{dt}$ $=xy-\beta z,$
where the canonical parameters are set to $\sigma=10$, $\rho=28$, and $\beta=8/3$. Under these settings, the system exhibits chaotic dynamics on a strange attractor in three-dimensional phase space.
The objective is to approximate the solution operator $\mathcal{G}\colon x_{0}\mapsto\mathbf{x}(t)$, where $\mathbf{x}(t)=(x(t),y(t),z(t))$ denotes the full trajectory over the time interval $t\in[0,3]$, given a scalar initial condition $x_{0}\in[10,15]$, with $y_{0}=z_{0}=12$ held fixed. This configuration reduces the input to a one-dimensional function-to-function mapping, while preserving the rich, sensitive behavior of the output trajectory.
A dataset of 10,00010{,}000 trajectories is generated using the classic Runge–Kutta method (RK45), each evaluated at 1000 uniformly spaced time steps. The input to the DeepONet branch network is the scalar $x_{0}$, and the trunk network is queried at time coordinates $t\in[0,3]$. The goal is to accurately recover the full state trajectory, including lobe transitions and fine-scale oscillations characteristic of chaotic dynamics.
Figure 8: Reconstructed Lorenz–63 attractor for a randomnly sampled initial condition
Figure 8 shows the reconstructed attractor for a representative test sample. While DeepONet departs from the ground truth early in the rollout, FEDONet maintains geometric coherence and captures both wings of the attractor. This result demonstrates the stabilizing effect of spectral enrichment on trajectory learning.
Figure 9: Phase-plane projection of the xxzz states. The vanilla DeepONet trajectory diverges from the expected orbit, showing geometric drift and incorrect lobe transitions. FEDONet reconstructs the spiral arms and inter-lobe dynamics with higher accuracy, preserving the overall topology.
In Figure 9, we examine the xxzz projection. The DeepONet prediction deviates from the true phase-space trajectory and fails to capture the correct lobe transitions. In contrast, FEDONet successfully recovers the attractor’s phase geometry, indicating more robust learning of long-term dependencies.
Figure 10: Pointwise prediction error in $x(t)$, $y(t)$, and $z(t)$ for a representative sample. The vanilla DeepONet exhibits compounding errors, especially after $t>1.5$, reflecting the accumulation of approximation drift. FEDONet achieves significantly lower errors and better error containment beyond lobe transitions, highlighting improved stability under chaotic evolution.
Figure 10 presents the pointwise prediction errors over time. DeepONet accumulates noticeable errors, particularly in the $x(t)$ and $z(t)$ components, which undergo rapid transitions between lobes. These deviations grow sharply after $t>1.5$, illustrating compounding approximation errors. FEDONet exhibits substantially lower and smoother error profiles, with better robustness during nonlinear state transitions.
Figure 11: Cumulative relative error over time. DeepONet exhibits rapid exponential error growth beyond $t>1.5$, characteristic of chaotic divergence. FEDONet accumulates error more gradually, indicating improved long-term stability and mitigation of runaway prediction drift.
In Figure 11, we observe that the cumulative error of DeepONet grows exponentially, characteristic of chaotic divergence. In contrast, FEDONet maintains a flatter cumulative error curve, demonstrating enhanced resilience against long-term instability and better preservation of trajectory fidelity.
Figure 12: Spectral energy density of predicted $x(t)$ trajectories compared with the ground truth. DeepONet underestimates energy in mid-to-high frequency bands, resulting in oversmoothing and loss of chaotic detail. FEDONet preserves the spectral tail and matches the broadband structure more accurately, contributing to improved phase and amplitude resolution.
Finally, Figure 12 compares the spectral content of the predicted and reference trajectories. The vanilla DeepONet fails to capture high-frequency modes, resulting in a smoothed trajectory that diverges from the correct dynamics. FEDONet, equipped with randomized Fourier features, preserves the broadband energy spectrum more accurately—an essential factor in recovering the fine-scale structure and temporal richness of chaotic systems.
These results collectively demonstrate that FEDONet offers improved phase alignment, lower pointwise and cumulative error, and superior spectral accuracy in chaotic dynamical systems. The introduction of Fourier embeddings into the trunk network acts as a spectral preconditioner, enabling the model to overcome inductive biases and improve fidelity under sensitivity to initial conditions.
3.4 Eikonal Equation
We consider the two-dimensional Eikonal equation, which governs the propagation of wavefronts at unit speed and arises in a wide range of applications including geometric optics, level-set methods, and computational geometry. The equation is given by:
$\|\nabla s(\mathbf{x})\|_{2}$ $=1,$ (17)
$s(\mathbf{x})$ $=0,\quad\mathbf{x}\in\partial\Omega,$
where $\mathbf{x}=(x,y)\in\mathbb{R}^{2}$ denotes the spatial coordinates, $\Omega\subset\mathbb{R}^{2}$ is the computational domain, and $\partial\Omega$ is its boundary. The solution $s(\mathbf{x})$ represents the signed distance from a point $\mathbf{x}\in\Omega$ to the boundary $\partial\Omega$, defined as:
[…]
with the distance function $d$ defined by:
$d(\mathbf{x},\partial\Omega):=\inf_{\mathbf{y}\in\partial\Omega}\|\mathbf{x}-\mathbf{y}\|_{2}.$ (19)
In this study, we aim to learn the solution operator that maps a binary mask of a two-dimensional geometry to its corresponding signed distance field (SDF). As a testbed, we construct a dataset of NACA 4-digit airfoil shapes, widely used in aerodynamics, and compute their SDFs within a structured computational domain.
A NACA 4-digit airfoil is parameterized by its maximum camber $m$, the location of maximum camber $p$, and the maximum thickness $t$. The mean camber line is given by:
$y_{c}(x)=\begin{cases}\frac{m}{p^{2}}(2px-x^{2}),&0\leq x<p,\\ \frac{m}{(1-p)^{2}}((1-2p)+2px-x^{2}),&p\leq x\leq 1,\end{cases}$ (20)
and its derivative is:
$\frac{dy_{c}}{dx}=\begin{cases}\frac{2m}{p^{2}}(p-x),&0\leq x<p,\\ \frac{2m}{(1-p)^{2}}(p-x),&p\leq x\leq 1.\end{cases}$ (21)
The thickness distribution $y_{t}(x)$ is defined as:
$y_{t}(x)=5t\left(0.2969\sqrt{x}-0.1260x-0.3516x^{2}+0.2843x^{3}-0.1036x^{4}\right).$ (22)
The upper and lower surface coordinates of the airfoil are computed as:
$x_{u}=x-y_{t}\sin(\theta),\quad$ $y_{u}=y_{c}+y_{t}\cos(\theta),$ (23)
$x_{l}=x+y_{t}\sin(\theta),\quad$ $y_{l}=y_{c}-y_{t}\cos(\theta),$ (24)
where $\theta=\tan^{-1}\left(\frac{dy_{c}}{dx}\right)$ is the local inclination angle of the camber line.
Each airfoil is embedded in a $256\times 256$ pixel computational grid and scaled to fit within a central subdomain. A binary mask is created to define the geometry, and the signed distance function is computed using the Euclidean Distance Transform (EDT) as:
$s(x,y)=d_{\text{out}}(x,y)-d_{\text{in}}(x,y),$ (25)
where $d_{\text{out}}$ and $d_{\text{in}}$ represent the unsigned distances outside and inside the airfoil boundary, respectively. To ensure numerical stability, the resulting SDF is normalized to the range [1,1][-1,1]:
$s_{\text{norm}}(x,y)=\frac{s(x,y)}{\max|s(x,y)|}.$ (26)
A total of 1250 distinct airfoil shapes are synthesized by randomly sampling parameters $m\in[0.01,0.09]$, $p\in[0.1,0.7]$, and $t\in[0.1,0.4]$. This dataset captures a wide range of geometrical variations and curvature profiles, providing a rigorous benchmark for learning PDE solution operators from geometric input.
Figure 13: Comparison of predicted signed distance functions (SDFs) for a representative airfoil geometry. Relative $L^{2}$ error: $2.728\times 10^{-2}$ for DeepONet, $0.912\times 10^{-2}$ for FEDONet.
As shown in Figure 13, FEDONet achieves more accurate recovery of the level-set structure. The DeepONet underperforms in regions of high curvature, particularly near the airfoil’s leading and trailing edges, where geometric non-smoothness introduces steep spatial gradients. FEDONet’s use of Fourier embeddings in the trunk network enhances its ability to resolve these features by expanding the function space with oscillatory basis functions aligned with the data’s spectral demands.
These results confirm that FEDONet offers improved resolution of high-gradient and discontinuous regions in PDE fields, even in geometry-conditioned tasks such as the Eikonal equation. This highlights its potential as a general-purpose surrogate for PDEs defined on complex, spatially varying domains.
3.5 Lorenz-96 Chaotic System
To evaluate the performance of our DeepONet-based neural operator models, we utilize the Lorenz-96 (L96) system—a well-known chaotic dynamical system commonly employed for benchmarking data-driven models and surrogate learning techniques in climate modeling and nonlinear dynamics. The L96 system is defined by the set of coupled ordinary differential equations:
$\frac{dx_{i}}{dt}=(x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F,\quad i=1,2,\dots,N$ (27)
with periodic boundary conditions: $x_{i-N}=x_{i}=x_{i+N}$. Here, $x_{i}$ denotes the state of the $i$-th variable, $N$ is the number of state variables, and $F$ is a constant forcing parameter that controls the nonlinearity and chaotic behavior of the system. For $F>8$, the system is known to exhibit strong chaos. In our experiments, we select $F=4.0$, which yields nonlinear but non-chaotic dynamics.
The L96 system is integrated forward in time using the classical fourth-order Runge-Kutta (RK4) method. Given an initial condition $\mathbf{x}_{0}\in\mathbb{R}^{N}$, the RK4 method computes the solution at the next time step using:
$\mathbf{x}_{n+1}=\mathbf{x}_{n}+\frac{\Delta t}{6}(k_{1}+2k_{2}+2k_{3}+k_{4})$ (28)
where the intermediate slopes $k_{1},k_{2},k_{3},k_{4}$ are computed using the Lorenz-96 time derivative evaluated at successive RK stages, and $\Delta t=0.01$ is the time step.
To construct a comprehensive dataset for training and evaluating neural operator models, we generate an ensemble of 10,00010{,}000 independent trajectories of the L96 system:
  • The system dimension is set to $N=40$.
  • Initial conditions are generated by perturbing the equilibrium state $x_{i}=F$ with small Gaussian noise:
    $\mathbf{x}_{0}=F\cdot\mathbf{1}+\epsilon\cdot\mathcal{N}(0,I),\quad\epsilon=10^{-3}$
  • Each trajectory is simulated for a total time of 15 seconds.
  • To remove transient behavior and capture steady-state dynamics, the first 10 seconds (burn-in period) are discarded.
  • The remaining 5 seconds of simulation are retained, corresponding to $500+1=501$ time steps with $\Delta t=0.01$.
The resulting dataset has shape (10,000,501,40)(10{,}000,501,40), where each sample corresponds to a trajectory of 40-dimensional state vectors over 501 time steps. This dataset captures the spatiotemporal evolution of the system from a diverse set of initial conditions and serves as a challenging testbed for learning nonlinear operator mappings.
Figure 14: Complete spatio-temporal field reconstruction comparison between DeepONet and FEDONet.
Figure 14 displays two representative hold-out trajectories of the L96 system. For each sample we plot the initial state, the reference spatio-temporal evolution, predictions from the DeepONet and FEDONet, together with the corresponding absolute-error fields.
  • Phase propagation. Both networks reproduce the diagonal stripe pattern that characterises L96 wave propagation, preserving phase speed and wavelength over the 500-step horizon. No visible drift or amplitude blow-up appears in either solution.
  • Error magnitude. The residual maps reveal only modest differences: in Sample 1 the RMS error of the vanilla model is $\mathcal{O}(10^{-1})$ while FEDONet trims it to the lower end of the same decade; Sample 2 shows the opposite ordering at isolated time–state patches. Overall, FEDONet yields a slightly lower global $\ell_{2}$ error but the gap is far narrower than that observed for the Poisson or Burgers benchmarks.
  • Interpretation. Because the L96 solution manifold is dominated by a single, slowly varying spatio-temporal mode, a shallow fully connected trunk already provides a sufficiently rich basis. Injecting random Fourier features therefore offers limited extra benefit, consistent with the quantitative numbers in Table 1, where both models attain sub-25 % relative error.
3.6 Allen–Cahn Equation
We evaluate the performance of the proposed Fourier-Embedded DeepONet (FEDONet) in modeling the one-dimensional Allen–Cahn equation—a nonlinear reaction-diffusion system widely used to describe phase separation processes in multi-component systems. The governing partial differential equation is given by:
$\frac{\partial u}{\partial t}=\epsilon\frac{\partial^{2}u}{\partial x^{2}}-5u^{3}+5u,\quad x\in[-1,1],\quad t\in[0,1],$ (29)
where $u(x,t)$ is the phase field variable and $\epsilon=10^{-4}$ denotes the diffusion coefficient. Periodic boundary conditions are enforced on both $u$ and its spatial derivative $\partial u/\partial x$, consistent with the physical assumption of a closed domain.
To construct a challenging dataset with rich spatial features, we generate 10,000 solution trajectories using an explicit Euler solver. The spatial and temporal grids are discretized using $\Delta x=0.01$ and $\Delta t=0.005$, yielding solution arrays of dimension $200\times 200$. Initial conditions are defined as:
$s(x)=\sum_{k=1}^{3}\left[a_{k}x^{2k}\cos(k\pi x)+b_{k}x^{2k}\sin(k\pi x)\right],$ (30)
where $a_{k},b_{k}\sim\mathcal{U}(0,1)$ are independent samples from the uniform distribution. This construction introduces multiscale structure and localized steep gradients into the initial data, posing a nontrivial test for surrogate learning.
Each training sample comprises a pair $(s(x),u(x,t))$, where $s(x)\in\mathbb{R}^{200\times 1}$ denotes the initial profile, and $u(x,t)\in\mathbb{R}^{200\times 200}$ is the full spatiotemporal solution. The goal is to approximate the solution operator $\mathcal{G}\colon s(x)\mapsto u(x,t)$, comparing the performance of vanilla DeepONet and FEDONet.
Figure 15: Best-performing test case for the Allen–Cahn equation. Relative $\ell_{2}$ error: $4.85\%$ for DeepONet, $2.27\%$ for FEDONet.
Figure 15 shows qualitative predictions for the best-performing test case. FEDONet sharply resolves the interface zones and phase transitions, while DeepONet produces smoother approximations with notable blurring near discontinuities. The spectral inductive bias introduced by randomized Fourier features enables FEDONet to retain high-frequency details crucial to phase-field dynamics.
Figure 16: Energy Spectrum and Pointwise absolute error across space and time for a representative sample.
The pointwise absolute error plot in Figure 16 emphasizes the localized improvements of FEDONet in capturing steep interfaces. DeepONet accumulates substantial errors in regions with high spatial gradients, whereas FEDONet maintains predictive accuracy throughout the domain, suggesting better handling of dissipative nonlinearities.
Figure 16 compares the spatial energy spectra of predicted and true solutions. FEDONet accurately tracks the ground truth spectrum across all wavenumbers, preserving the high-frequency tail associated with sharp phase boundaries. In contrast, DeepONet suffers from spectral attenuation at $k>20$, revealing an inability to fully capture dissipative and oscillatory dynamics inherent in the Allen–Cahn system.
These results collectively demonstrate that the incorporation of spectral priors via Fourier embeddings in FEDONet enhances both pointwise accuracy and spectral fidelity. The improved resolution of phase interfaces and better retention of high-frequency information position FEDONet as a robust surrogate model for nonlinear, multiscale PDEs such as the Allen–Cahn equation.
3.7 Kuramoto–Sivashinsky Equation
We evaluate the performance of FEDONet on the one-dimensional Kuramoto–Sivashinsky (KS) equation, a canonical model that exhibits spatiotemporal chaos due to the combined effects of nonlinear advection, destabilizing diffusion, and stabilizing fourth-order hyperviscosity. The governing equation is given by:
$\begin{split}&\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}+\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{4}u}{\partial x^{4}}=0,\quad x\in[0,L],\;t\in[0,T],\\ &u(x,0)=u_{0}(x),\quad x\in[0,L],\end{split}$ (31)
where $u(x,t)$ is the scalar field of interest, and $L=24$, $T=50$ denote the spatial and temporal domains respectively. The initial condition $u_{0}(x)$ is synthesized as a randomized Fourier series:
$u_{0}(x)=\sum_{n=1}^{4}C_{n}\sin\left(\frac{nx}{L}\right),\quad C_{n}\sim\mathcal{N}(0,1),$ (32)
which introduces low-frequency coherent structures that quickly evolve into chaotic waveforms under the KS dynamics.
We generate 10,000 such initial conditions and simulate the corresponding solution fields using a pseudo-spectral solver. The objective is to learn the operator $\mathcal{G}\colon u_{0}(x)\mapsto u(x,t)$ using both the standard DeepONet and the proposed FEDONet models. This problem serves as a stringent benchmark due to the sensitivity of chaotic systems to initial conditions and their broadband frequency content.
Figure 17: Reconstruction of a median-performing KS trajectory. Relative $L^{2}$ error: FEDONet = 16.36%, DeepONet = 89.63%.
Figure 17 shows the reconstructed spatiotemporal field for a median test case. The FEDONet model effectively captures the nonlinear evolution of the solution, preserving the wave envelope, phase dynamics, and intermittent modulation typical of KS attractors. In stark contrast, the vanilla DeepONet produces incoherent and smoothed reconstructions, with pronounced phase lag and energy collapse beyond $t>20$. This degradation is reflected in a high relative error of $89.63\%$ compared to just $16.36\%$ for FEDONet.
In Figure 18, we highlight the best-performing test case. FEDONet achieves a relative $L^{2}$ error of only $10.77\%$, accurately modeling both the fine-scale chaotic filaments and large-scale rolling structures. Notably, the trajectory remains consistent with the ground truth over long integration times, demonstrating that the spectral trunk can stabilize predictions and suppress drift. Such performance underscores the model’s ability to retain energy in high-frequency modes—essential for modeling non-periodic fine-grained behavior that typifies chaotic fields.
Figure 18: Best-performing KS test sample. FEDONet retains the amplitude and frequency content of the chaotic trajectory, yielding a low relative $L^{2}$ error of $10.77\%$.
Figure 19: Worst-performing test sample. Even in failure cases, FEDONet remains qualitatively reasonable. Relative $L^{2}$ error: FEDONet = $58.02\%$, versus near-complete divergence for DeepONet.
Figure 19 displays the worst-case test result. While the FEDONet prediction exhibits growing deviations from the reference field, it still retains discernible wave structures and phase coherence, with an error of $58.02\%$. In comparison, the vanilla DeepONet fails entirely, producing numerical artifacts, amplitude distortion, and full trajectory collapse. This robustness under extreme conditions emphasizes the value of frequency-enhanced representations in learning dynamical system operators.
Figure 20 plots the training loss curves for both models. FEDONet converges faster and exhibits more stable optimization behavior, attributable to the improved conditioning of the input space via randomized Fourier features. This aligns with observations from kernel learning and feature engineering literature, where frequency-based transformations are known to facilitate smoother descent landscapes and gradient propagation.
Figure 20: Training convergence comparison. FEDONet achieves faster and more stable optimization due to implicit spectral regularization.
Figure 21: Distribution of relative $L^{2}$ errors across 512 unseen test samples. FEDONet maintains lower error in both mean and variance, demonstrating improved robustness and generalization in chaotic regimes.
Finally, Figure 21 summarizes performance statistics across 512 randomly selected test cases. FEDONet consistently achieves lower relative $L^{2}$ errors with reduced variance, indicating better generalization and resilience to initial condition perturbations.
In totality, the Kuramoto–Sivashinsky benchmark clearly illustrates the limitations of standard DeepONet trunks in handling chaotic, broadband dynamics. The introduction of Fourier embeddings into the trunk network provides an effective inductive bias that promotes spectral alignment, temporal coherence, and long-horizon accuracy. FEDONet emerges as a compelling model for operator learning in complex regimes, particularly those characterized by sensitive dependence on initial conditions and multiscale feature interactions.
4 Summary
In this work, we propose the Fourier-Embedded DeepONet (FEDONet) as a principled extension to the classical DeepONet architecture for learning nonlinear operators governed by partial differential equations (PDEs). By injecting randomized Fourier features into the trunk input, FEDONet expands the hypothesis space of DeepONets, offering enhanced spectral representation, improved numerical conditioning, and faster convergence during training—all without increasing runtime complexity. FEDONet consistently outperforms the vanilla DeepONet. Notably, it delivers dramatic reductions in relative $\ell_{2}$ error—most prominently in multiscale and high-frequency scenarios such as the Kuramoto–Sivashinsky equation (76% $\rightarrow$ 21%).
Table 1: Relative $\ell_{2}$ error (%) for each benchmark.
Dataset DeepONet FEDONet
2-D Poisson (elliptic) 5.80 2.32
1-D Burgers (conv./diff.) 4.01 2.63
Lorenz–63 (chaotic ODE) 2.03 0.46
Lorenz–96 (chaotic ODE) 23.40 22.90
Allen–Cahn (phase–field) 10.61 5.50
Kuramoto–Sivashinsky (spatio-chaos) 75.11 20.62
Eikonal (Hamilton–Jacobi) 2.15 1.12
Table 2: Qualitative performance across seven benchmark problems.
Method 2-D Poisson 1-D Burgers Lorenz–63 Lorenz–96 Allen–Cahn KS Eikonal
DeepONet
FEDONet ✓✓ ✓✓ ✓✓ ✓✓
  • Legend: ✓ decent;  ✓✓ excellent;  ✗ sub-par.
Beyond pointwise accuracy, FEDONet preserves physical consistency across modalities. Energy spectrum analyses reveal that FEDONet better captures inertial and dissipation range behavior, avoiding the artificial spectral pile-up observed in DeepONets. In geometry-sensitive tasks like the Eikonal equation, it reconstructs sharp features and discontinuities with higher fidelity.
Taken together, these results establish FEDONet as a robust, scalable, and architecture-agnostic enhancement for neural operator learning. The proposed spectral trunk serves as a drop-in replacement that aligns the inductive biases of DeepONets with the spectral demands of complex PDE systems, providing a promising path forward for efficient surrogate modeling in scientific machine learning.
5 Future Work
The Fourier-Embedded DeepONet (FEDONet) architecture presents a compelling foundation for enhancing operator learning, yet several promising extensions remain unexplored. A natural next step is to generalize beyond sinusoidal embeddings by investigating alternative basis functions in the trunk network. Wavelets, Chebyshev polynomials, or orthogonal polynomials tailored to specific PDE families could offer improved localization or boundary adaptivity, especially for systems exhibiting sharp interfaces or multi-scale dynamics. Additionally, hybridizing spectral embeddings with structured architectures such as graph-based trunks or encoder–decoder networks may further boost performance on irregular domains or hierarchical solution spaces.
Another avenue involves making the Fourier embedding parameters learnable. While the current design uses fixed random frequency matrices, enabling adaptive tuning of these frequencies could allow the model to better align with the intrinsic spectral content of the data. This modification could yield a more expressive and data-efficient hypothesis space while preserving the training stability benefits of approximate whitening. Finally, incorporating uncertainty quantification and physics-informed constraints into the FEDONet framework—via Bayesian formulations or PINO-style residual losses—may further enhance robustness, interpretability, and generalization in scientific computing tasks involving sparse or noisy data.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
This work was supported in part by the AFOSR Grant FA9550-24-1-0327.
Data availability
Data supporting the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Whitening Effect of Fourier Feature Embeddings
Let $\phi(\zeta)=\sqrt{2}\,[\sin(2\pi B\zeta),\cos(2\pi B\zeta)]\in\mathbb{R}^{2M}$ be the Fourier feature embedding of an input $\zeta\in\mathbb{R}^{d}$, where each row of the matrix $B\in\mathbb{R}^{M\times d}$ is sampled i.i.d. from a Gaussian distribution:
$B_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (33)
Assume further that the input $\zeta$ is uniformly distributed over a compact domain, e.g., $\zeta\sim\mathcal{U}([0,1]^{d})$. Under these assumptions, we show that the random Fourier features $\phi(\zeta)$ are approximately whitened, i.e.,
$\mathbb{E}_{\zeta}[\phi(\zeta)\phi(\zeta)^{\top}]\approx I_{2M}.$ (34)
To see this, consider a single frequency row vector $B_{i}\in\mathbb{R}^{1\times d}$ drawn from $B$. The corresponding $2$-dimensional block of the embedding is:
$\phi_{i}(\zeta)=\sqrt{2}\begin{bmatrix}\sin(2\pi B_{i}\cdot\zeta)\\ \cos(2\pi B_{i}\cdot\zeta)\end{bmatrix}$. (35)
We now compute the second-order statistics of this embedding block under the assumption that $B_{i}\cdot\zeta$ is uniformly distributed over $[0,1]$ (a reasonable approximation due to the randomness in $B_{i}$ and the uniformity of $\zeta$). Using standard trigonometric integrals, we obtain:
$\mathbb{E}_{\zeta}[\sin^{2}(2\pi B_{i}\cdot\zeta)]$ $\approx\frac{1}{2},$ (36)
$\mathbb{E}_{\zeta}[\cos^{2}(2\pi B_{i}\cdot\zeta)]$ $\approx\frac{1}{2},$ (37)
$\mathbb{E}_{\zeta}[\sin(2\pi B_{i}\cdot\zeta)\cos(2\pi B_{i}\cdot\zeta)]$ $\approx 0.$ (38)
Therefore, the unscaled covariance matrix for this block is approximately:
$\mathbb{E}_{\zeta}\left[\begin{bmatrix}\sin(2\pi B_{i}\cdot\zeta)\\ \cos(2\pi B_{i}\cdot\zeta)\end{bmatrix}\begin{bmatrix}\sin(2\pi B_{i}\cdot\zeta)&\cos(2\pi B_{i}\cdot\zeta)\end{bmatrix}\right]\approx\frac{1}{2}I_{2}.$ (39)
Scaling by the $\sqrt{2}$ factor yields:
$\mathbb{E}_{\zeta}[\phi_{i}(\zeta)\phi_{i}(\zeta)^{\top}]\approx I_{2}.$ (40)
Since the rows of $B$ are sampled independently and the embeddings for each $B_{i}$ are uncorrelated, the full embedding vector $\phi(\zeta)\in\mathbb{R}^{2M}$ satisfies:
$\mathbb{E}_{\zeta}[\phi(\zeta)\phi(\zeta)^{\top}]\approx I_{2M}.$ (41)
Hence, the random Fourier features are approximately whitened. This whitening effect implies that the feature coordinates are uncorrelated and have unit variance. In practice, this leads to improved numerical conditioning and convergence in gradient-based optimization, particularly via the Neural Tangent Kernel (NTK), which benefits from isotropic input distributions. As such, Fourier embeddings not only enhance expressivity but also promote well-conditioned training dynamics.
Appendix B Theoretical Justification for Fourier Embedding Superset Property
In this appendix, we formally justify the claim stated in Section 2.1 that the hypothesis space induced by Fourier embeddings strictly contains the vanilla hypothesis space of coordinate-input networks. We present this as a lemma and provide constructive proof using classical function approximation theory.
Let $\mathcal{H}_{\text{vanilla}}$ be the hypothesis space of functions representable by a feedforward neural network with input $\zeta\in\mathbb{R}^{d}$, and let $\mathcal{H}_{\phi}$ be the hypothesis space of functions representable by a similar network where the input is first mapped via the Fourier feature embedding:
$\phi(\zeta)=\left[\sin(2\pi B\zeta),\cos(2\pi B\zeta)\right],\quad B\in\mathbb{R}^{M\times d},\;B_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (42)
Then, under mild regularity assumptions (e.g., compact input domain $K\subset\mathbb{R}^{d}$ and continuous activation functions), we have:
$\mathcal{H}_{\text{vanilla}}\subsetneq\mathcal{H}_{\phi}.$ (43)
We divide the proof into two parts.
Part I: Inclusion. Let $f\in\mathcal{H}_{\text{vanilla}}$ be a function representable by a neural network of the form:
$f(\zeta)=\sum_{j=1}^{N}a_{j}\,\sigma(w_{j}^{\top}\zeta+b_{j}).$ (44)
Since each $\zeta_{i}$ (coordinate function) can be approximated by a Fourier series on a compact domain $K$, and the Fourier embedding $\phi(\zeta)$ contains these sinusoidal components, it follows that a function of $\zeta$ can be emulated by a function of $\phi(\zeta)$. In particular, for every $f\in\mathcal{H}_{\text{vanilla}}$ and $\varepsilon>0$, there exists $\tilde{f}\in\mathcal{H}_{\phi}$ such that:
$\sup_{\zeta\in K}|f(\zeta)-\tilde{f}(\zeta)|<\varepsilon.$ (45)
Therefore, $\mathcal{H}_{\text{vanilla}}\subseteq\mathcal{H}_{\phi}$.
Part II: Strict Containment. Consider the function $f(\zeta)=\sin(100\pi\zeta)$ over $\zeta\in[0,1]$. Due to the spectral bias of neural networks [49], this high-frequency function is difficult to approximate using shallow networks on raw input. However, if the frequency matrix $B$ used in the embedding $\phi(\zeta)$ includes rows close to $100$, then:
$\sin(2\pi B_{j}\zeta)\approx\sin(100\pi\zeta),$
making $f$ easily representable in $\mathcal{H}_{\phi}$.
Since $f(\zeta)$ cannot be efficiently approximated in $\mathcal{H}_{\text{vanilla}}$ (requiring significantly more parameters or depth), but is trivially accessible in $\mathcal{H}_{\phi}$, we conclude that the inclusion is strict:
$\mathcal{H}_{\text{vanilla}}\subsetneq\mathcal{H}_{\phi}.$ (46)
This result formalizes the intuition that Fourier embeddings expand the expressivity of the network and can be interpreted as a form of spectral preconditioning for operator learning models such as DeepONets.
Appendix C Spectral Interpretation of Fourier-Embedded DeepONets
The use of Fourier embeddings in the trunk network induces a natural spectral decomposition in the DeepONet output. Specifically, the composition of the trunk and branch networks can be interpreted as a data-driven spectral projection:
$\mathcal{G}(u)(\zeta)\approx\sum_{k=1}^{p}\langle u,\varphi_{k}\rangle\,\psi_{k}(\zeta),$ (47)
where $\{\varphi_{k}\}_{k=1}^{p}$ are data-dependent basis functions learned from the branch network, and $\{\psi_{k}(\zeta)\}_{k=1}^{p}$ are basis functions defined by the trunk network, shaped by the sinusoidal feature embedding $\phi(\zeta)$.
When the trunk network is fed sinusoidal embeddings of the form
$\phi(\zeta)=[\cos(2\pi\omega_{1}^{\top}\zeta),\dots,\cos(2\pi\omega_{M}^{\top}\zeta),\sin(2\pi\omega_{1}^{\top}\zeta),\dots,\sin(2\pi\omega_{M}^{\top}\zeta)],$
the basis functions $\psi_{k}(\zeta)$ are learned combinations of trigonometric modes. This allows the model to represent oscillatory and localized patterns that are characteristic of multiscale physical systems.
The representation in Equation (47) mirrors classical Galerkin methods, where the solution is projected onto a known orthonormal basis:
$u(x,t)\approx\sum_{n=1}^{N}a_{n}(t)\,\phi_{n}(x),$ (48)
with the coefficients $a_{n}(t)$ governed by PDE dynamics. In DeepONet, both the basis functions $\psi_{k}$ and the projection coefficients $\langle u,\varphi_{k}\rangle$ are learned directly from data, enabling generalization beyond handcrafted bases.
Unlike classical global bases such as Fourier or Chebyshev functions, the neural trunk network can generate localized, adaptive basis functions. This enhances the model’s ability to approximate functions with sharp transitions, discontinuities, or compactly supported features, which are common in real-world PDE systems. We may formalize the DeepONet architecture as a composition of a projection and synthesis operator:
$\Pi_{\varphi}(u):=[\langle u,\varphi_{1}\rangle,\dots,\langle u,\varphi_{p}\rangle]^{\top},\quad\mathcal{S}_{\psi}(\zeta):=[\psi_{1}(\zeta),\dots,\psi_{p}(\zeta)]^{\top},$
yielding the output:
$\mathcal{G}(u)(\zeta)\approx\mathcal{S}_{\psi}(\zeta)^{\top}\Pi_{\varphi}(u),$
which resembles a spectral synthesis operation of the form:
$f(x)\approx\sum_{k}\hat{f}_{k}\phi_{k}(x).$
This interpretation leads to several insights:
  1. 1.
    Fourier-embedded DeepONets naturally align with the dominant spectral modes of the target operator, facilitating more efficient approximation of high-frequency content.
  2. 2.
    The learned basis can adapt to complex patterns in the data, unlike fixed basis expansions.
  3. 3.
    The operator learning framework parallels well-established numerical discretization techniques, such as Petrov–Galerkin and spectral methods, but operates in a fully data-driven regime.
In summary, Fourier embeddings serve as a principled mechanism for spectral lifting, enabling DeepONet to perform a learned Galerkin-style decomposition of operators. This perspective bridges classical PDE approximation theory with modern deep learning, offering both interpretability and flexibility in learning complex function-to-function mappings.

### v2

@@
 Abstract
-Deep Operator Networks (DeepONets) have recently emerged as powerful data-driven frameworks for learning nonlinear operators, particularly suited for approximating solutions to partial differential equations (PDEs).
-Despite their promising capabilities, the standard implementation of DeepONets, which typically employs fully connected linear layers in the trunk network, can encounter limitations in capturing complex spatial structures inherent to various PDEs.
-To address this, we introduce Fourier-embedded trunk networks within the DeepONet architecture, leveraging random Fourier feature mappings to enrich spatial representation capabilities.
-Our proposed Fourier-embedded DeepONet, FEDONet demonstrates superior performance compared to the traditional DeepONet across a comprehensive suite of PDE-driven datasets, including the two-dimensional Poisson equation, Burgers’ equation, the Lorenz-63 chaotic system, Eikonal equation, Allen–Cahn equation, Kuramoto–Sivashinsky equation, and the Lorenz-96 system.
-Empirical evaluations of FEDONet consistently show significant improvements in solution reconstruction accuracy, with average relative $\ell_{2}$ performance gains ranging between 2-3x compared to the DeepONet baseline.
-This study highlights the effectiveness of Fourier embeddings in enhancing neural operator learning, offering a robust and broadly applicable methodology for PDE surrogate modeling.
+[…]
 1 Introduction
-Partial differential equations (PDEs) underpin models ranging from turbulent-flow dynamics and heat diffusion to electromagnetics and biological transport.
-Classical numerical solvers—finite-difference, finite-element and spectral methods—provide high-fidelity solutions but incur a prohibitive computational burden when faced with high-dimensional, multiscale or stiff systems [1, 2, 3].
-The quest for rapid surrogates has a long pedigree: universal-approximation theorems for neural networks [4, 5, 6, 7], reduced-order techniques based on proper-orthogonal decomposition (POD) [8, 9], kernel–Gaussian-process surrogates [10], and mesh-free radial-basis schemes [11, 12].
-Yet, translating these ideas into *function-to-function* (operator) mappings that respect arbitrary discretisations remained elusive.
-The modern wave of operator learning addresses this gap by training neural networks that act directly on infinite-dimensional function spaces and generalise across grid resolutions.
-Deep Operator Networks (DeepONets) [13] and the broader Neural Operator family [14, 15, 16] establish a branch–trunk architecture with rigorous operator–universal-approximation guarantees.
-Fourier Neural Operators (FNOs) extend this paradigm through global spectral convolutions, delivering impressive speed-ups on parametric PDEs [17].
-Subsequent variants exploit multiwavelets [18, 19], convolutional kernels [20], multi-grid tensorisation [21, 22] and geometry-aware deformations [23] to boost resolution or handle complex domains.
-Beyond kernel choices, researchers have tackled discontinuities [24], derivative supervision [25], irregular meshes via graph trunks [26], active multi-resolution sampling [27], and U-Net/wavelet hybrids [28, 29].
-Random Fourier features have proved especially effective at encoding high-frequency structure in standard deep nets [30], yet remain under-explored inside operator networks whose trunk is typically a plain fully connected multilayer perceptron (MLP).
-To incorporate governing laws and uncertainty, Physics-Informed Neural Networks (PINNs) [31, 32] evolved into Physics-Informed Neural Operators (PINO) [33, 34] and their variational or pseudo-physics relatives [35, 36].
-Complementary efforts embed operators in latent or invertible spaces for data efficiency and bidirectional inference [37, 38, 39, 40].
-Transformer-style attention has recently been adopted for operator learning [41, 42, 43, 44, 45] and for geometry generalisation [46].
-Such architectures underpin landmark climate-scale surrogates such as FourCastNet [47] and its amortised FNO successor [48].
-Despite the rich taxonomy above, the trunk of conventional DeepONets, charged with encoding spatial/temporal coordinates, remains a shallow MLP.
-This design struggles to capture the oscillatory or multi-scale patterns ubiquitous in nonlinear PDE solutions, leading to sub-optimal sample efficiency and accuracy.
-Recent trunk alternatives (e.g.
-graph [26] or wavelet U-Net [28]) introduce architectural complexity or domain-specific preprocessing.
-We propose a Fourier-Embedded DeepONet (FEDONet) that replaces the raw coordinates fed into the trunk with a compact random Fourier-feature map [30].
-The embedding supplies an implicit spectrum of sinusoidal basis functions, endowing the trunk with a richer hypothesis space at negligible parameter cost.
-FEDONet preserves the original branch–trunk separation, is plug-and-play with existing DeepONet codebases, and adds no runtime overhead beyond a single matrix multiplication.
-  1. (i)
-    We formulate the Fourier embedding of trunk coordinates, derive its effect on the operator neural tangent kernel, and show that it strictly enlarges the approximation class of DeepONets.
-  2. (ii)
-    We deliver a thorough benchmark across seven PDE families—steady Poisson, viscous Burgers, Allen–Cahn, Kuramoto–Sivashinsky, Eikonal, Lorenz–63 and Lorenz–96—demonstrating 48.93% average reduction in relative $\ell^{2}$ error versus a tuned vanilla DeepONet.
-  3. (iii)
-    We analyse hyperparameter sensitivity (mapping size and scale) and empirically confirm that the Fourier trunk improves convergence speed without sacrificing discretisation invariance.
-By injecting a principled spectral bias into the DeepONet trunk, FEDONet closes the performance gap to frequency-domain operators such as FNO while retaining the flexibility and locality advantages of the original architecture.
+Partial differential equations underpin models ranging from turbulent‐flow dynamics and heat diffusion to electromagnetics and biological transport.
+Classical numerical solvers such as finite‐difference, finite‐element and spectral methods provide high‐fidelity solutions but incur a prohibitive computational burden when faced with high‐dimensional, multiscale or stiff systems [1, 2, 3].
+Efforts to construct efficient surrogate models have a long history, spanning the universal approximation theorems for neural networks[4, 5, 6, 7], reduced‐order techniques based on proper‐orthogonal decomposition [8, 9], kernel Gaussian process surrogates [10], and mesh‐free radial‐basis schemes [11, 12].
+However, extending these approaches to learn function‐to‐function mappings capable of generalizing across discretizations, boundary conditions, and input configurations has remained a long standing challenge.
+The recent emergence of operator learning addresses this long-standing challenge by training neural networks that act directly on infinite-dimensional function spaces, enabling generalization across discretizations and grid resolutions.
+Deep Operator Networks [13] pioneered the branch–trunk architecture for learning nonlinear operators, while subsequent Neural Operator formulations [14, 15, 16] extended these ideas with rigorous operator–approximation guarantees and scalable architectures.
+Building on these foundations, Fourier Neural Operators [17] employ global spectral convolutions to capture nonlocal correlations efficiently, achieving remarkable speed-ups and accuracy gains in parametric PDE learning.
+Subsequent developments have further advanced this paradigm through multiwavelet representations [18, 19], convolutional kernels [20], multigrid tensorizations [21, 22], and geometry-aware spectral deformations [23], collectively enhancing scalability, resolution fidelity, and geometric adaptability.
+Beyond architectural and kernel innovations, recent research has expanded operator learning along multiple frontiers.
+Efforts addressing solution irregularities have focused on handling discontinuities [24], incorporating derivative supervision for enhanced gradient fidelity [25], and enabling learning on irregular meshes through graph-based trunk networks [26].
+Adaptive and multi-resolution strategies have been introduced to improve sampling efficiency [27], while hybrid encoder-decoder frameworks leveraging U-Net and wavelet representations further enhance multi-scale expressivity [28, 29].
+To embed physical constraints and quantify uncertainty, the paradigm of Physics-Informed Neural Networks (PINNs) [30, 31] has been generalized into Physics-Informed Neural Operators [32, 33] and related variational or pseudo-physics extensions [34, 35].
+Parallel efforts have explored latent and invertible operator formulations to achieve data-efficient and bidirectional inference [36, 37, 38, 39].
+Transformer based attention mechanisms have also been adapted for operator learning [40, 41, 42, 43, 44] and for robust geometry generalization [45].
+Together, these advances have enabled high-fidelity, climate-scale surrogates such as FourCastNet [46] and its amortized Fourier Neural Operator successor [47], underscoring the maturity and scalability of the operator-learning paradigm.
+Despite the recent operator-learning architectures, the trunk network of conventional DeepONets remains a shallow multilayer perceptron.
+This design often struggles to represent the oscillatory and multi-scale structures that characterize nonlinear PDE solutions, resulting in reduced sample efficiency and reconstruction accuracy.
+While recent alternatives such as graph-based [26] or wavelet-enhanced U-Net trunks [28] offer improved expressivity, they introduce significant architectural complexity and domain-specific preprocessing requirements.
+In contrast, random Fourier features have proven remarkably effective at encoding high-frequency information in standard deep networks [48], yet their potential within operator-learning frameworks remains largely unexplored, where the trunk is typically implemented as a simple fully connected MLP.
+To address this gap, we propose the Fourier-Embedded Deep Operator Network (FEDONet), which augments the DeepONet trunk by replacing raw spatial coordinates with a compact random Fourier-feature embedding [48].
+This embedding implicitly introduces a spectrum of sinusoidal basis functions, enriching the trunk’s representational capacity with negligible additional parameters.
+FEDONet retains the original branch–trunk separation, integrates seamlessly with existing DeepONet implementations, and incurs no runtime overhead beyond a single matrix multiplication.
+We formally introduce the Fourier embedding of trunk coordinates and analyze its theoretical implications through the lens of the operator neural tangent kernel, demonstrating that the embedding strictly enlarges the approximation class of standard DeepONets.
+A comprehensive benchmark is carried spanning seven canonical PDE families - steady Poisson, viscous Burgers, Allen–Cahn, Kuramoto–Sivashinsky, Eikonal, Lorenz–63, and Lorenz–96.
+By embedding a principled spectral bias into the trunk network, FEDONet bridges the performance gap to frequency-domain operators, while preserving the architectural simplicity, flexibility, and locality benefits of the original DeepONet formulation.
+Section 2 presents the formulation of the Fourier-embedded Deep Operator Network (FEDONet), detailing the embedding construction, its theoretical connection to the operator neural tangent kernel, and its implications for approximation capacity.
+Section 3 describes the benchmark setup and implementation details across the seven PDE families considered.
+Section 4 reports quantitative and spectral analyses of FEDONet performance, including convergence behavior, and comparisons with baseline operator models.
+Finally, Section 5 summarizes the main findings and outlines directions for future research in spectral operator learning and physics-constrained network design.
 2 Methodology
@@
 2.1 Fourier Embeddings as Spectral Preconditioners
-Standard MLPs exhibit a well-known spectral bias — they learn low-frequency components faster than high-frequency ones [49].
+Standard MLPs exhibit a well-known spectral bias - they learn low-frequency components faster than high-frequency ones [49].
 This bottleneck is particularly harmful in operator learning problems involving sharp gradients, multiscale dynamics, or oscillatory fields.
-To address this, we propose injecting fixed Fourier embeddings into the trunk network input. Given a coordinate $\zeta\in\mathbb{R}^{d}$, we define a randomized Fourier feature map:
+To address this, we propose injecting fixed Fourier embeddings into the trunk network input illustrated in Fig.
+1.
+Given a coordinate $\zeta\in\mathbb{R}^{d}$, we define a randomized Fourier feature map:
 $\phi(\zeta)=\left[\sin(2\pi B\zeta),\cos(2\pi B\zeta)\right],\quad B_{ij}\sim\mathcal{N}(0,\sigma^{2}),$ (4)
-where $B\in\mathbb{R}^{M\times d}$ is a Gaussian frequency matrix. Instead of feeding raw coordinates to the trunk, we input $\phi(\zeta)$ — enriching its spectral content.
-The DeepONet prediction then becomes:
+where $B\in\mathbb{R}^{M\times d}$ is a Gaussian frequency matrix. Instead of feeding raw coordinates to the trunk, we input $\phi(\zeta)$ - enriching its spectral content.
+The FEDONet prediction then becomes:
 $\mathcal{G}_{\theta}(u)(\zeta)=B_{\theta}(u)\cdot T_{\theta}(\phi(\zeta))=\sum_{k=1}^{p}b_{k}(u)\cdot t_{k}(\phi(\zeta)).$ (5)
@@
 effectively lifting $\zeta$ into a high-dimensional basis of sinusoidal functions.
-By Bochner’s theorem [50], such embeddings approximate shift-invariant kernels — implicitly kernelizing the trunk network.
+By Bochner’s theorem [50], such embeddings approximate shift-invariant kernels - implicitly kernelizing the trunk network.
 This expands its representational capacity, enabling better learning of high-frequency targets.
@@
 We adopt a supervised learning framework to train the proposed FEDONet model.
-The objective is to approximate a target operator $G$ by minimizing the prediction error over a dataset of paired input–output functions.
+The objective is to approximate a target operator $G$ by minimizing the prediction error over a dataset of paired input-output functions.
 Each dataset element consists of a function $u^{(i)}\in\mathcal{U}$ and corresponding evaluations of the output $s^{(i)}_{j}=G(u^{(i)})(\zeta^{(i)}_{j})$ at query locations $\zeta^{(i)}_{j}\in\mathcal{Y}$.
@@
 Learning rate schedules are selected based on the problem characteristics to ensure stable and efficient convergence.
-Each mini-batch consists of randomly sampled function–query pairs $(u,\zeta)$, enabling scalable and efficient training across diverse domains.
+Each mini-batch consists of randomly sampled function-query pairs $(u,\zeta)$, enabling scalable and efficient training across diverse domains.
 Notably, the Fourier embedding layer $\phi(\zeta)$ is fixed and non-trainable throughout training.
@@
 3.1 2D Poisson Equation
-We created a synthetic dataset comprising numerical solutions of the two-dimensional (2D) Poisson equation, a fundamental PDE commonly employed in modeling steady-state processes such as heat conduction, electrostatics, and diffusion phenomena.
-The mathematical formulation of the Poisson equation is given by:
+We constructed a synthetic dataset comprising numerical solutions to the canonical two-dimensional Poisson equation. The governing equation is
 $\nabla^{2}u(x,y)=f(x,y),\quad(x,y)\in[0,1]^{2},$ (9)
-accompanied by homogeneous Dirichlet boundary conditions:
-$u(x,y)\big{|}_{\partial\Omega}=0.$ (10)
-In this context, $u(x,y)$ denotes the solution field, while $f(x,y)$ represents a specified source term or coefficient.
-To generate the dataset, we synthesized $N=10,000$ distinct instances of the coefficient function $f$ over a uniform spatial grid of resolution $128\times 128$.
-Each instance of the coefficient field was produced using Gaussian Random Fields (GRFs) characterized by specific spectral parameters, notably $\alpha=3$ and $\tau=3$, resulting in:
-$f\sim\text{GRF}(\alpha=3,\tau=3).$ (11)
-Once these coefficient fields were created, we numerically solved the associated Poisson equation to obtain the corresponding solution fields $u(x,y)$.
-The numerical discretization employed a finite difference approach, utilizing a conventional five-point stencil approximation of the Laplacian operator.
-By discretizing the spatial domain $\Omega=[0,1]^{2}$ into a $128\times 128$ structured grid, the PDE was transformed into a linear algebraic system:
-$A\mathbf{u}=b,$ (12)
-where $\mathbf{u}\in\mathbb{R}^{16384}$ denotes the vectorized solution field and $b\in\mathbb{R}^{16384}$ is the vector representation of the source term $f$.
-The discretized Laplacian matrix $A$ is defined explicitly as:
-[…]
-with homogeneous Dirichlet boundary conditions enforced directly by setting boundary rows of $A$ as identity matrices and corresponding entries in $f$ to zero.
-Ultimately, this methodology yielded 10,000 independent pairs of coefficient-source fields and their respective solutions, forming a robust dataset suitable for training, validating, and benchmarking neural operator models.
-The dataset facilitates a comprehensive evaluation of the trade-offs in computational efficiency, accuracy, and reconstruction performance of our proposed neural operator architectures.
-Figure 2 presents a representative test sample from the two–dimensional Poisson dataset.
-The input forcing field is shown on the left, followed by the exact solution, the prediction of a vanilla DeepONet, and the corresponding point-wise error.
-The bottom row repeats the comparison for our Fourier-Embedded DeepONet (FEDONet).
-  • Spatial reconstruction.
-FEDONet reproduces the lobed extrema, smooth gradients and corner behaviour of the ground truth almost indistinguishably, whereas the vanilla network exhibits visible blurring and slight spatial shifts of the high-magnitude regions.
-The residual of the vanilla model oscillates between $\pm 1.5\times 10^{-3}$, displaying high-frequency artefacts; FEDONet reduces this range to $\pm 3\times 10^{-4}$, a five-fold improvement that manifests as a low-amplitude, coherent pattern.
-  • Energy spectra.
-Figure 3 plots the angle-integrated energy spectrum $E_{k}$ for four randomly chosen hold-out fields (samples 29, 15, 24 and 17).
-Both networks match the reference spectrum at large scales ($k\lesssim 5$), but qualitative differences arise in the inertial and dissipation ranges:
-    1. (a)
-      The vanilla DeepONet (purple) consistently over-estimates energy for $k\gtrsim 8$, producing a spectral “pile-up” that grows by nearly an order of magnitude in sample 24—mirroring the spurious small-scale oscillations seen in its error map.
-    2. (b)
-      FEDONet (red) tracks the ground truth (dark blue) almost perfectly up to the Nyquist limit, deviating by at most a factor of two in the far dissipation tail and thereby preserving the correct energy cascade.
-Figure 2: Model reconstruction comparison for the 2D Poisson Dataset.
-Together, the spatial and spectral diagnostics confirm that Fourier embeddings in the trunk endow DeepONets with the high-frequency resolution and stability required for accurate operator learning: FEDONet not only slashes point-wise errors but also eliminates the artificial energy plateau that plagues fully-connected trunks.
+subject to homogeneous Dirichlet boundary conditions,
+$u(x,y)\big|_{\partial\Omega}=0,$ (10)
+where $u(x,y)$ denotes the potential field and $f(x,y)$ represents a prescribed source distribution.
+A total of N=10, [^0] N=10{,}000 distinct source fields were synthesized using Gaussian random fields (GRFs) with spectral parameters $\alpha=3$ and $\tau=3$,
+$f\sim\text{GRF}(\alpha=3,\tau=3),$ (11)
+on a uniform $128\times 128$ grid.
+For each realization of $f$, the corresponding solution $u$ was obtained by numerically solving the Poisson equation using a standard five-point finite-difference discretization of the Laplacian operator.
+The spatial domain $\Omega=[0,1]^{2}$ was thus transformed into a structured algebraic system,
+$A\mathbf{u}=\mathbf{b},$ (12)
+where $\mathbf{u},\mathbf{b}\in\mathbb{R}^{16384}$ denote the vectorized forms of the solution and source fields, respectively.
+The coefficient matrix $A$ was defined using a conventional five-point stencil with Dirichlet conditions enforced by setting boundary rows to the identity and zeroing the corresponding entries in $\mathbf{b}$.
+This process yielded 10,000 pairs of input-output fields suitable for training, validation, and testing of neural operator architectures.
+Figure 2: Comparison of predicted and reference solution fields for the two-dimensional Poisson equation.
+Figure 2 shows a representative test case from the Poisson dataset.
+The top row depicts the input forcing, reference solution, prediction from the baseline DeepONet, and the corresponding pointwise error, while the bottom row presents results from the proposed Fourier-Embedded DeepONet (FEDONet).
+FEDONet reproduces the reference field with sharper gradients and more accurate spatial localization of extrema, whereas the baseline model exhibits noticeable blurring and displacement of high-magnitude regions.
+The residual field of the vanilla DeepONet fluctuates within $\pm 1.5\times 10^{-3}$ and displays fine-scale artifacts, while FEDONet reduces the range to $\pm 3\times 10^{-4}$ and yields a smoother, spatially coherent error distribution, corresponding to an approximate fivefold improvement in reconstruction accuracy.
+Spectral accuracy was assessed using the angle-integrated energy spectrum $E_{k}$ (Fig.
+3).
+Both models reproduce the reference spectrum at large scales ($k\lesssim 5$), but differences arise in the intermediate and dissipative ranges.
+The baseline DeepONet consistently overestimates energy for $k\gtrsim 8$, resulting in a spurious buildup of high-wavenumber content that mirrors the oscillatory residuals seen in physical space.
+In contrast, FEDONet closely follows the reference spectrum across the entire range, deviating by less than a factor of two near the Nyquist limit and maintaining a physically consistent energy decay.
 Figure 3: Energy spectrum comparison between the ground truth, vanilla and Fourier DeepOnet reconstructions for the 2D Poisson Dataset.
+Taken together, the spatial and spectral analyses demonstrate that the Fourier embedding enhances the trunk network’s ability to capture high-frequency features and improves overall stability.
+FEDONet achieves lower pointwise errors while eliminating the artificial spectral plateau that characterizes conventional fully connected trunks.
 3.2 Burgers’ Equation
@@
 FEDONet consistently outperforms DeepONet, with a mean error reduction from 4.01% to 2.63%, and a narrower spread in error distribution indicating increased reliability.
-Figure 4: Burgers Relative $\ell_{2}$ error across 128 unseen samples for DeepONet and FEDONet.
 Figure 5 presents the best-performing test sample, showing the ground truth solution alongside predictions from both models.
@@
 In contrast, FEDONet produces significantly sharper and more accurate reconstructions, faithfully capturing shock fronts and fine-scale dynamics.
-Figure 5: Best-performing test sample: Ground truth vs. DeepONet vs. FEDONet prediction.
+Figure 4: Burgers Relative $\ell_{2}$ error across 128 unseen samples for DeepONet and FEDONet.
 The energy spectra of the predicted and ground truth fields, shown in Figure 6, further emphasize this advantage.
@@
 This demonstrates that the Fourier embeddings effectively bridge the model’s spectral bias, enabling it to better represent high-frequency dynamics essential to Burgers-type flows.
+Figure 5: Best-performing test sample for Burgers’ Equation: Ground truth vs. DeepONet vs. FEDONet prediction.
 These results underscore the importance of aligning model inductive bias with the physical spectrum of the target PDE.
 The enhanced performance of FEDONet in both physical and spectral domains highlights its potential as a more robust and generalizable operator learner for convection-dominated systems.
-Figure 6: Temporal Evolution of Pointwise Error and Energy Spectrum: Best-Performing Sample vs. Ground Truth
+Figure 6: Temporal Evolution of Pointwise Error and Energy Spectrum for Burgers’ Equation
 Figure 7 shows the convergence of training losses for different batch sizes.
 FEDONet consistently converges faster and to lower loss values across all training configurations, indicating improved optimization characteristics and stability during training.
-Figure 7: Training loss convergence for the 1D Burgers’ Equation: Comparison of Vanilla and Fourier Embedded DeepONets trained with varying batch sizes.
+Figure 7: Training loss convergence for the 1D Burgers’ Equation - Comparison of Vanilla and Fourier Embedded DeepONets trained with varying batch sizes.
 3.3 Lorenz–63 System
@@
 This configuration reduces the input to a one-dimensional function-to-function mapping, while preserving the rich, sensitive behavior of the output trajectory.
-A dataset of 10,00010{,}000 trajectories is generated using the classic Runge–Kutta method (RK45), each evaluated at 1000 uniformly spaced time steps.
+A dataset of 10,00010{,}000 trajectories is generated using the classic Runge-Kutta method (RK45), each evaluated at 1000 uniformly spaced time steps.
 The input to the DeepONet branch network is the scalar $x_{0}$, and the trunk network is queried at time coordinates $t\in[0,3]$.
@@
 The vanilla DeepONet fails to capture high-frequency modes, resulting in a smoothed trajectory that diverges from the correct dynamics.
-FEDONet, equipped with randomized Fourier features, preserves the broadband energy spectrum more accurately—an essential factor in recovering the fine-scale structure and temporal richness of chaotic systems.
+FEDONet, equipped with randomized Fourier features, preserves the broadband energy spectrum more accurately - an essential factor in recovering the fine-scale structure and temporal richness of chaotic systems.
 These results collectively demonstrate that FEDONet offers improved phase alignment, lower pointwise and cumulative error, and superior spectral accuracy in chaotic dynamical systems.
@@
 This dataset captures a wide range of geometrical variations and curvature profiles, providing a rigorous benchmark for learning PDE solution operators from geometric input.
-Figure 13: Comparison of predicted signed distance functions (SDFs) for a representative airfoil geometry.
+Figure 13: Comparison of predicted signed distance functions (SDFs) for a representative airfoil geometry for Eikonal Equation.
 Relative $L^{2}$ error: $2.728\times 10^{-2}$ for DeepONet, $0.912\times 10^{-2}$ for FEDONet.
@@
 3.5 Lorenz-96 Chaotic System
-To evaluate the performance of our DeepONet-based neural operator models, we utilize the Lorenz-96 (L96) system—a well-known chaotic dynamical system commonly employed for benchmarking data-driven models and surrogate learning techniques in climate modeling and nonlinear dynamics.
-The L96 system is defined by the set of coupled ordinary differential equations:
-$\frac{dx_{i}}{dt}=(x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F,\quad i=1,2,\dots,N$ (27)
-with periodic boundary conditions: $x_{i-N}=x_{i}=x_{i+N}$.
-Here, $x_{i}$ denotes the state of the $i$-th variable, $N$ is the number of state variables, and $F$ is a constant forcing parameter that controls the nonlinearity and chaotic behavior of the system.
-For $F>8$, the system is known to exhibit strong chaos.
-In our experiments, we select $F=4.0$, which yields nonlinear but non-chaotic dynamics.
-The L96 system is integrated forward in time using the classical fourth-order Runge-Kutta (RK4) method.
-Given an initial condition $\mathbf{x}_{0}\in\mathbb{R}^{N}$, the RK4 method computes the solution at the next time step using:
-$\mathbf{x}_{n+1}=\mathbf{x}_{n}+\frac{\Delta t}{6}(k_{1}+2k_{2}+2k_{3}+k_{4})$ (28)
-where the intermediate slopes $k_{1},k_{2},k_{3},k_{4}$ are computed using the Lorenz-96 time derivative evaluated at successive RK stages, and $\Delta t=0.01$ is the time step.
-To construct a comprehensive dataset for training and evaluating neural operator models, we generate an ensemble of 10,00010{,}000 independent trajectories of the L96 system:
-  • The system dimension is set to $N=40$.
-  • Initial conditions are generated by perturbing the equilibrium state $x_{i}=F$ with small Gaussian noise:
-    $\mathbf{x}_{0}=F\cdot\mathbf{1}+\epsilon\cdot\mathcal{N}(0,I),\quad\epsilon=10^{-3}$
-  • Each trajectory is simulated for a total time of 15 seconds.
-  • To remove transient behavior and capture steady-state dynamics, the first 10 seconds (burn-in period) are discarded.
-  • The remaining 5 seconds of simulation are retained, corresponding to $500+1=501$ time steps with $\Delta t=0.01$.
-The resulting dataset has shape (10,000,501,40)(10{,}000,501,40), where each sample corresponds to a trajectory of 40-dimensional state vectors over 501 time steps.
-This dataset captures the spatiotemporal evolution of the system from a diverse set of initial conditions and serves as a challenging testbed for learning nonlinear operator mappings.
-Figure 14: Complete spatio-temporal field reconstruction comparison between DeepONet and FEDONet.
-Figure 14 displays two representative hold-out trajectories of the L96 system.
-For each sample we plot the initial state, the reference spatio-temporal evolution, predictions from the DeepONet and FEDONet, together with the corresponding absolute-error fields.
-  • Phase propagation.
-Both networks reproduce the diagonal stripe pattern that characterises L96 wave propagation, preserving phase speed and wavelength over the 500-step horizon.
-No visible drift or amplitude blow-up appears in either solution.
-  • Error magnitude.
-The residual maps reveal only modest differences: in Sample 1 the RMS error of the vanilla model is $\mathcal{O}(10^{-1})$ while FEDONet trims it to the lower end of the same decade; Sample 2 shows the opposite ordering at isolated time–state patches.
-Overall, FEDONet yields a slightly lower global $\ell_{2}$ error but the gap is far narrower than that observed for the Poisson or Burgers benchmarks.
-  • Interpretation.
-Because the L96 solution manifold is dominated by a single, slowly varying spatio-temporal mode, a shallow fully connected trunk already provides a sufficiently rich basis.
-Injecting random Fourier features therefore offers limited extra benefit, consistent with the quantitative numbers in Table 1, where both models attain sub-25 % relative error.
+To assess model performance on nonlinear dynamical systems, we consider the Lorenz-96 (L96) model, a prototypical chaotic system widely used for testing data-driven surrogates in climate modeling and nonlinear forecasting.
+The governing equations are
+$\frac{dx_{i}}{dt}=(x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F,\quad i=1,2,\dots,N,$ (26)
+subject to periodic boundary conditions $x_{i-N}=x_{i}=x_{i+N}$.
+Here, $x_{i}$ denotes the state of the $i$-th variable, $N$ is the system dimension, and $F$ is a constant forcing parameter that determines the level of nonlinearity and chaos.
+For $F>8$, the system exhibits strong chaotic behavior; in this study, we set $F=4.0$ to obtain moderately nonlinear yet stable dynamics.
+The system is integrated using a classical fourth-order Runge–Kutta (RK4) scheme with a fixed time step $\Delta t=0.01$:
+$\mathbf{x}_{n+1}=\mathbf{x}_{n}+\frac{\Delta t}{6}(k_{1}+2k_{2}+2k_{3}+k_{4}),$ (27)
+where $k_{1},k_{2},k_{3},k_{4}$ are the intermediate RK stages computed from the Lorenz–96 derivatives.
+A dataset of 10,00010{,}000 independent trajectories is generated for training and evaluation. Each trajectory is initialized as
+$\mathbf{x}_{0}=F\cdot\mathbf{1}+\epsilon\,\mathcal{N}(0,I),\quad\epsilon=10^{-3},$
+with $N=40$ and integrated for a total physical time of 15 s.
+The initial 10 s are discarded as a transient, and the remaining 5 s (501 time steps) are retained to capture quasi-stationary behavior.
+The resulting dataset has dimensions (10,000,501,40)(10{,}000,501,40), representing 10,000 spatio-temporal trajectories of 40-dimensional state vectors.
+Figure 14: Lorenz-96 - Complete spatio-temporal field reconstruction comparison between DeepONet and FEDONet.
+Figure 14 shows representative test trajectories from the L96 dataset.
+Each panel depicts the initial condition, reference evolution, predictions from the baseline DeepONet and the proposed Fourier-Embedded DeepONet (FEDONet), and their respective absolute-error fields.
+Both models reproduce the characteristic diagonal stripe patterns associated with wave-like phase propagation, maintaining amplitude and wavelength stability over the full 500-step forecast horizon.
+No visible drift or blow-up is observed in either solution.
+Quantitatively, the residual fields indicate comparable performance.
+In the first trajectory, the vanilla DeepONet attains a root-mean-square (RMS) error on the order of $\mathcal{O}(10^{-1})$, while FEDONet slightly reduces this magnitude; in the second trajectory, the ordering reverses locally, but the overall difference remains small.
+On average, FEDONet achieves marginally lower global $\ell_{2}$ errors, though the improvement is less pronounced than in the Poisson or Burgers benchmarks.
+This outcome reflects the inherent structure of the L96 dynamics: the solution manifold is dominated by a slowly varying, low-dimensional spatio-temporal mode, which a shallow fully connected trunk can already approximate effectively.
+Consequently, introducing random Fourier features provides limited additional benefit, consistent with the quantitative results in Table 1, where both models achieve sub-25% relative error.
 3.6 Allen–Cahn Equation
-We evaluate the performance of the proposed Fourier-Embedded DeepONet (FEDONet) in modeling the one-dimensional Allen–Cahn equation—a nonlinear reaction-diffusion system widely used to describe phase separation processes in multi-component systems.
+We evaluate the performance of the proposed Fourier-Embedded DeepONet (FEDONet) in modeling the one-dimensional Allen–Cahn equation nonlinear reaction-diffusion system widely used to describe phase separation processes in multi-component systems.
 The governing partial differential equation is given by:
@@
 The spectral inductive bias introduced by randomized Fourier features enables FEDONet to retain high-frequency details crucial to phase-field dynamics.
-Figure 16: Energy Spectrum and Pointwise absolute error across space and time for a representative sample.
 The pointwise absolute error plot in Figure 16 emphasizes the localized improvements of FEDONet in capturing steep interfaces.
 DeepONet accumulates substantial errors in regions with high spatial gradients, whereas FEDONet maintains predictive accuracy throughout the domain, suggesting better handling of dissipative nonlinearities.
+Figure 16: Allen-Cahn - Energy Spectrum and Pointwise absolute error across space and time for a representative sample.
 Figure 16 compares the spatial energy spectra of predicted and true solutions.
@@
 This problem serves as a stringent benchmark due to the sensitivity of chaotic systems to initial conditions and their broadband frequency content.
-Figure 17: Reconstruction of a median-performing KS trajectory. Relative $L^{2}$ error: FEDONet = 16.36%, DeepONet = 89.63%.
+Figure 17: Median-performing KS trajectory. Relative $L^{2}$ error: FEDONet = 16.36%, DeepONet = 89.63%.
 Figure 17 shows the reconstructed spatiotemporal field for a median test case.
 The FEDONet model effectively captures the nonlinear evolution of the solution, preserving the wave envelope, phase dynamics, and intermittent modulation typical of KS attractors.
-In stark contrast, the vanilla DeepONet produces incoherent and smoothed reconstructions, with pronounced phase lag and energy collapse beyond $t>20$.
+In stark contrast, the vanilla DeepONet produces incoherent and smoothed reconstructions, with pronounced phase lag and energy collapse.
 This degradation is reflected in a high relative error of $89.63\%$ compared to just $16.36\%$ for FEDONet.
@@
 Notably, the trajectory remains consistent with the ground truth over long integration times, demonstrating that the spectral trunk can stabilize predictions and suppress drift.
-Such performance underscores the model’s ability to retain energy in high-frequency modes—essential for modeling non-periodic fine-grained behavior that typifies chaotic fields.
+Such performance underscores the model’s ability to retain energy in high-frequency modes essential for modeling non-periodic fine-grained behavior that typifies chaotic fields.
 Figure 18: Best-performing KS test sample. FEDONet retains the amplitude and frequency content of the chaotic trajectory, yielding a low relative $L^{2}$ error of $10.77\%$.
-Figure 19: Worst-performing test sample. Even in failure cases, FEDONet remains qualitatively reasonable. Relative $L^{2}$ error: FEDONet = $58.02\%$, versus near-complete divergence for DeepONet.
+Figure 19: Worst-performing KS test sample. Even in failure cases, FEDONet remains qualitatively reasonable. Relative $L^{2}$ error: FEDONet = $58.02\%$, versus near-complete divergence for DeepONet.
 Figure 19 displays the worst-case test result.
@@
 This aligns with observations from kernel learning and feature engineering literature, where frequency-based transformations are known to facilitate smoother descent landscapes and gradient propagation.
-Figure 20: Training convergence comparison. FEDONet achieves faster and more stable optimization due to implicit spectral regularization.
-Figure 21: Distribution of relative $L^{2}$ errors across 512 unseen test samples.
+Figure 20: KS Training convergence comparison. FEDONet achieves faster and more stable optimization due to implicit spectral regularization.
+Figure 21: Distribution of relative $L^{2}$ errors across 512 unseen test samples for KS Equation.
 FEDONet maintains lower error in both mean and variance, demonstrating improved robustness and generalization in chaotic regimes.
@@
 FEDONet consistently achieves lower relative $L^{2}$ errors with reduced variance, indicating better generalization and resilience to initial condition perturbations.
-In totality, the Kuramoto–Sivashinsky benchmark clearly illustrates the limitations of standard DeepONet trunks in handling chaotic, broadband dynamics.
+In totality, the Kuramoto-Sivashinsky benchmark clearly illustrates the limitations of standard DeepONet trunks in handling chaotic, broadband dynamics.
 The introduction of Fourier embeddings into the trunk network provides an effective inductive bias that promotes spectral alignment, temporal coherence, and long-horizon accuracy.
@@
 In this work, we propose the Fourier-Embedded DeepONet (FEDONet) as a principled extension to the classical DeepONet architecture for learning nonlinear operators governed by partial differential equations (PDEs).
-By injecting randomized Fourier features into the trunk input, FEDONet expands the hypothesis space of DeepONets, offering enhanced spectral representation, improved numerical conditioning, and faster convergence during training—all without increasing runtime complexity.
+By injecting randomized Fourier features into the trunk input, FEDONet expands the hypothesis space of DeepONets, offering enhanced spectral representation, improved numerical conditioning, and faster convergence during training all without increasing runtime complexity.
 FEDONet consistently outperforms the vanilla DeepONet.
-Notably, it delivers dramatic reductions in relative $\ell_{2}$ error—most prominently in multiscale and high-frequency scenarios such as the Kuramoto–Sivashinsky equation (76% $\rightarrow$ 21%).
+Notably, it delivers dramatic reductions in relative $\ell_{2}$ error, most prominently in multiscale and high-frequency scenarios such as the Kuramoto-Sivashinsky equation (76% $\rightarrow$ 21%).
 Table 1: Relative $\ell_{2}$ error (%) for each benchmark.
@@
 Eikonal (Hamilton–Jacobi) 2.15 1.12
-Table 2: Qualitative performance across seven benchmark problems.
-Method 2-D Poisson 1-D Burgers Lorenz–63 Lorenz–96 Allen–Cahn KS Eikonal
-DeepONet
-FEDONet ✓✓ ✓✓ ✓✓ ✓✓
-  • Legend: ✓ decent;  ✓✓ excellent;  ✗ sub-par.
 Beyond pointwise accuracy, FEDONet preserves physical consistency across modalities.
@@
 This modification could yield a more expressive and data-efficient hypothesis space while preserving the training stability benefits of approximate whitening.
-Finally, incorporating uncertainty quantification and physics-informed constraints into the FEDONet framework—via Bayesian formulations or PINO-style residual losses—may further enhance robustness, interpretability, and generalization in scientific computing tasks involving sparse or noisy data.
+Finally, incorporating uncertainty quantification and physics-informed constraints into the FEDONet framework via Bayesian formulations or PINO style residual losses may further enhance robustness, interpretability, and generalization in scientific computing tasks involving sparse or noisy data.
 Declaration of competing interest

### v3

@@
 Abstract
+Deep Operator Networks (DeepONets) have recently emerged as powerful data-driven frameworks for learning nonlinear operators, particularly suited for approximating solutions to partial differential equations.
+Despite their promising capabilities, the standard implementation of DeepONets, which typically employs fully connected linear layers in the trunk network, can encounter limitations in capturing complex spatial structures inherent to various PDEs.
+To address this limitation, we introduce Fourier-Embedded trunk networks within the DeepONet architecture, leveraging random fourier feature mappings to enrich spatial representation capabilities.
+Our proposed Fourier-Embedded DeepONet, FEDONet demonstrates superior performance compared to the traditional DeepONet across a comprehensive suite of PDE-driven datasets, including the two-dimensional Poisson, Burgers’, Lorenz-63, Eikonal, Allen-Cahn, and the Kuramoto-Sivashinsky equation.
+FEDONet delivers consistently superior reconstruction accuracy across all benchmark PDEs, with particularly large relative $L^{2}$ error reductions observed in chaotic and stiff systems.
+This study highlights the effectiveness of Fourier embeddings in enhancing neural operator learning, offering a robust and broadly applicable methodology for PDE surrogate modeling.
+keywords:
+Neural Operator Learning , Fourier Feature Embeddings , Deep Operator Networks , Partial Differential Equations
+journal: Journal of Computational Physics
 […]
+organization=Mechanical and Aerospace Engineering, addressline=1512 Middle Drive, University of Tennessee, city=Knoxville, postcode=37996-2210, state=TN, country=USA
 1 Introduction
@@
 Efforts to construct efficient surrogate models have a long history, spanning the universal approximation theorems for neural networks[4, 5, 6, 7], reduced‐order techniques based on proper‐orthogonal decomposition [8, 9], kernel Gaussian process surrogates [10], and mesh‐free radial‐basis schemes [11, 12].
-However, extending these approaches to learn function‐to‐function mappings capable of generalizing across discretizations, boundary conditions, and input configurations has remained a long standing challenge.
+Early data‐driven approaches also sought to learn governing equations directly from trajectory data using deep neural networks [13] However, extending these approaches to learn function‐to‐function mappings capable of generalizing across discretizations, boundary conditions, and input configurations has remained a long standing challenge.
 The recent emergence of operator learning addresses this long-standing challenge by training neural networks that act directly on infinite-dimensional function spaces, enabling generalization across discretizations and grid resolutions.
-Deep Operator Networks [13] pioneered the branch–trunk architecture for learning nonlinear operators, while subsequent Neural Operator formulations [14, 15, 16] extended these ideas with rigorous operator–approximation guarantees and scalable architectures.
+Deep Operator Networks [14] pioneered the branch-trunk architecture for learning nonlinear operators, while subsequent Neural Operator formulations [15, 16, 17] extended these ideas with rigorous operator-approximation guarantees and scalable architectures.
 Building on these foundations, Fourier Neural Operators [18] employ global spectral convolutions to capture nonlocal correlations efficiently, achieving remarkable speed-ups and accuracy gains in parametric PDE learning.
+Related ideas have been explored in the context of learning PDE dynamics in modal or Fourier space [19], where the evolution operator is approximated directly on spectral coefficients rather than physical fields, highlighting the efficacy of frequency-space representations for capturing multiscale dynamics.
 Subsequent developments have further advanced this paradigm through multiwavelet representations [20, 21], convolutional kernels [22], multigrid tensorizations [23, 24], and geometry-aware spectral deformations [25], collectively enhancing scalability, resolution fidelity, and geometric adaptability.
@@
 This embedding implicitly introduces a spectrum of sinusoidal basis functions, enriching the trunk’s representational capacity with negligible additional parameters.
-FEDONet retains the original branch–trunk separation, integrates seamlessly with existing DeepONet implementations, and incurs no runtime overhead beyond a single matrix multiplication.
+FEDONet retains the original branch-trunk separation, integrates seamlessly with existing DeepONet implementations, and incurs no runtime overhead beyond a single matrix multiplication.
 We formally introduce the Fourier embedding of trunk coordinates and analyze its theoretical implications through the lens of the operator neural tangent kernel, demonstrating that the embedding strictly enlarges the approximation class of standard DeepONets.
-A comprehensive benchmark is carried spanning seven canonical PDE families - steady Poisson, viscous Burgers, Allen–Cahn, Kuramoto–Sivashinsky, Eikonal, Lorenz–63, and Lorenz–96.
+A comprehensive benchmark is carried spanning seven canonical PDE families - steady Poisson, viscous Burgers, Allen-Cahn, Kuramoto-Sivashinsky, Eikonal, and Lorenz-63.
 By embedding a principled spectral bias into the trunk network, FEDONet bridges the performance gap to frequency-domain operators, while preserving the architectural simplicity, flexibility, and locality benefits of the original DeepONet formulation.
-Section 2 presents the formulation of the Fourier-embedded Deep Operator Network (FEDONet), detailing the embedding construction, its theoretical connection to the operator neural tangent kernel, and its implications for approximation capacity.
-Section 3 describes the benchmark setup and implementation details across the seven PDE families considered.
-Section 4 reports quantitative and spectral analyses of FEDONet performance, including convergence behavior, and comparisons with baseline operator models.
-Finally, Section 5 summarizes the main findings and outlines directions for future research in spectral operator learning and physics-constrained network design.
+Section  2 presents the formulation of the Fourier-Embedded Deep Operator Network (FEDONet), detailing the embedding construction, its theoretical connection to the operator neural tangent kernel, and its implications for approximation capacity.
+Section  3 describes the benchmark setup and implementation details across the seven PDE families considered.
+Sections  4 and  5 summarize the main findings and outlines directions for future research in spectral operator learning and physics-constrained network design.
 2 Methodology
@@
 Given a coordinate $\zeta\in\mathbb{R}^{d}$, we define a randomized Fourier feature map:
-$\phi(\zeta)=\left[\sin(2\pi B\zeta),\cos(2\pi B\zeta)\right],\quad B_{ij}\sim\mathcal{N}(0,\sigma^{2}),$ (4)
-where $B\in\mathbb{R}^{M\times d}$ is a Gaussian frequency matrix. Instead of feeding raw coordinates to the trunk, we input $\phi(\zeta)$ - enriching its spectral content.
+$\phi(\zeta)=\left[\sin(2\pi Z\zeta),\cos(2\pi Z\zeta)\right],\quad Z_{ij}\sim\mathcal{N}(0,\sigma^{2}),$ (4)
+where $Z\in\mathbb{R}^{M\times d}$ is a Gaussian frequency matrix. Instead of feeding raw coordinates to the trunk, we input $\phi(\zeta)$ - enriching its spectral content.
 The FEDONet prediction then becomes:
@@
 This process yielded 10,000 pairs of input-output fields suitable for training, validation, and testing of neural operator architectures.
-Figure 2: Comparison of predicted and reference solution fields for the two-dimensional Poisson equation.
+Figure 2: Solving a 2D Poisson equation: Comparison of predicted and reference solution fields of DeepONet and FEDONet
 Figure 2 shows a representative test case from the Poisson dataset.
-The top row depicts the input forcing, reference solution, prediction from the baseline DeepONet, and the corresponding pointwise error, while the bottom row presents results from the proposed Fourier-Embedded DeepONet (FEDONet).
-FEDONet reproduces the reference field with sharper gradients and more accurate spatial localization of extrema, whereas the baseline model exhibits noticeable blurring and displacement of high-magnitude regions.
-The residual field of the vanilla DeepONet fluctuates within $\pm 1.5\times 10^{-3}$ and displays fine-scale artifacts, while FEDONet reduces the range to $\pm 3\times 10^{-4}$ and yields a smoother, spatially coherent error distribution, corresponding to an approximate fivefold improvement in reconstruction accuracy.
-Spectral accuracy was assessed using the angle-integrated energy spectrum $E_{k}$ (Fig.
-3).
-Both models reproduce the reference spectrum at large scales ($k\lesssim 5$), but differences arise in the intermediate and dissipative ranges.
-The baseline DeepONet consistently overestimates energy for $k\gtrsim 8$, resulting in a spurious buildup of high-wavenumber content that mirrors the oscillatory residuals seen in physical space.
-In contrast, FEDONet closely follows the reference spectrum across the entire range, deviating by less than a factor of two near the Nyquist limit and maintaining a physically consistent energy decay.
+The top row depicts the input forcing field, exact solution, prediction from the baseline DeepONet, and the corresponding pointwise error, while the bottom row presents results from the proposed Fourier-Embedded DeepONet (FEDONet).
+Both models were evaluated on $1000$ randomly selected unseen test samples, yielding average relative $L^{2}$ errors of $1.41\%$ for the DeepONet and $1.09\%$ for the FEDONet.
+FEDONet reproduces the reference field with sharper gradients and more accurate spatial localization of extrema.
+The residual field of the vanilla DeepONet displays fine-scale artifacts, while FEDONet yields a smoother, spatially coherent error distribution, corresponding to an approximate fivefold improvement in reconstruction accuracy.
 Figure 3: Energy spectrum comparison between the ground truth, vanilla and Fourier DeepOnet reconstructions for the 2D Poisson Dataset.
+Spectral accuracy was assessed using the angle-integrated energy spectrum $E_{k}$ shown in Fig.
+3.
+Both models reproduce the reference spectrum at large scales ($k\lesssim 50$), but differences arise in the intermediate and dissipative ranges.
+The baseline DeepONet overestimates energy for $k\gtrsim 90$, resulting in a spurious buildup of high-wavenumber content that mirrors the oscillatory residuals seen in physical space.
+In contrast, FEDONet closely follows the reference spectrum across the entire range.
 Taken together, the spatial and spectral analyses demonstrate that the Fourier embedding enhances the trunk network’s ability to capture high-frequency features and improves overall stability.
@@
 3.2 Burgers’ Equation
-The Burgers’ equation is a well-known 1D viscous equation. It has been studied extensively for DeepONets in prior literature. The equation is given by:
+The Burgers’ equation is a well-known 1D viscous equation. It has been studied extensively for DeepONets in prior literature [53]. The equation is given by:
 $\frac{\partial u}{\partial t}(x,t)+u\frac{\partial u}{\partial x}(x,t)=\nu\frac{\partial^{2}u}{\partial x^{2}}(x,t),\>\>\>\>\>\forall(x,t)\in[0,1]\times[0,1]$ (13)
@@
 The objective is to learn the nonlinear operator $\mathcal{G}$ mapping $s(x)$ to the full spatio-temporal solution $u(x,t)$ using both vanilla DeepONet and the proposed FEDONet.
-Figure 4 shows the relative $\ell_{2}$ errors over 128 unseen test samples.
-FEDONet consistently outperforms DeepONet, with a mean error reduction from 4.01% to 2.63%, and a narrower spread in error distribution indicating increased reliability.
-Figure 5 presents the best-performing test sample, showing the ground truth solution alongside predictions from both models.
+Both models were evaluated on $250$ unseen test samples, yielding average relative $L^{2}$ errors of $6.43\%$ for the DeepONet and $4.86\%$ for the FEDONet.
+Figure 4 presents prediction for a test sample, showing the ground truth solution alongside predictions from both models.
 While DeepONet captures the coarse solution structure, it struggles with steep gradients and exhibits visible smoothing near shock regions.
 In contrast, FEDONet produces significantly sharper and more accurate reconstructions, faithfully capturing shock fronts and fine-scale dynamics.
-Figure 4: Burgers Relative $\ell_{2}$ error across 128 unseen samples for DeepONet and FEDONet.
+Figure 4: Solving a parametric Burgers’ equation: Comparison of predicted spatio-temporal field for a random initial condition. Relative $L^{2}$ error: $4.09\%$ for DeepONet, $1.49\%$ for FEDONet.
 The energy spectra of the predicted and ground truth fields, shown in Figure 5, further emphasize this advantage.
@@
 This demonstrates that the Fourier embeddings effectively bridge the model’s spectral bias, enabling it to better represent high-frequency dynamics essential to Burgers-type flows.
-Figure 5: Best-performing test sample for Burgers’ Equation: Ground truth vs. DeepONet vs. FEDONet prediction.
-These results underscore the importance of aligning model inductive bias with the physical spectrum of the target PDE.
 The enhanced performance of FEDONet in both physical and spectral domains highlights its potential as a more robust and generalizable operator learner for convection-dominated systems.
-Figure 6: Temporal Evolution of Pointwise Error and Energy Spectrum for Burgers’ Equation
+Figure 5: Energy spectrum and pointwise absolute error across space and time for a representative sample for Burgers’ Equation.
 Figure 6 shows the convergence of training losses for different batch sizes.
 FEDONet consistently converges faster and to lower loss values across all training configurations, indicating improved optimization characteristics and stability during training.
-Figure 7: Training loss convergence for the 1D Burgers’ Equation - Comparison of Vanilla and Fourier Embedded DeepONets trained with varying batch sizes.
-3.3 Lorenz–63 System
-The Lorenz–63 system is a widely studied chaotic dynamical system that serves as a benchmark for evaluating the long-term stability and predictive accuracy of surrogate models.
+Figure 6: Training loss convergence for the 1D Burgers’ Equation - Comparison of DeepONet and FEDONet trained with varying batch sizes.
+3.3 Lorenz-63 System
+The Lorenz-63 system is a widely studied chaotic dynamical system that serves as a benchmark for evaluating the long-term stability and predictive accuracy of surrogate models.
 It is governed by the following system of coupled nonlinear ordinary differential equations:
@@
 The goal is to accurately recover the full state trajectory, including lobe transitions and fine-scale oscillations characteristic of chaotic dynamics.
-Figure 8: Reconstructed Lorenz–63 attractor for a randomnly sampled initial condition
+Figure 7: Reconstructed Lorenz-63 attractor for a randomnly sampled initial condition for trained DeepONet and FEDONet
 Figure 7 shows the reconstructed attractor for a representative test sample.
@@
 Figure 8: Phase-plane projection of the xxzz states.
-The vanilla DeepONet trajectory diverges from the expected orbit, showing geometric drift and incorrect lobe transitions.
-FEDONet reconstructs the spiral arms and inter-lobe dynamics with higher accuracy, preserving the overall topology.
 In Figure 8, we examine the xxzz projection.
@@
 Figure 9: Pointwise prediction error in $x(t)$, $y(t)$, and $z(t)$ for a representative sample.
-The vanilla DeepONet exhibits compounding errors, especially after $t>1.5$, reflecting the accumulation of approximation drift.
-FEDONet achieves significantly lower errors and better error containment beyond lobe transitions, highlighting improved stability under chaotic evolution.
 Figure 9 presents the pointwise prediction errors over time.
@@
 Figure 10: Cumulative relative error over time.
-DeepONet exhibits rapid exponential error growth beyond $t>1.5$, characteristic of chaotic divergence.
-FEDONet accumulates error more gradually, indicating improved long-term stability and mitigation of runaway prediction drift.
 In Figure 10, we observe that the cumulative error of DeepONet grows exponentially, characteristic of chaotic divergence.
@@
 Figure 11: Spectral energy density of predicted $x(t)$ trajectories compared with the ground truth.
-DeepONet underestimates energy in mid-to-high frequency bands, resulting in oversmoothing and loss of chaotic detail.
-FEDONet preserves the spectral tail and matches the broadband structure more accurately, contributing to improved phase and amplitude resolution.
 Finally, Figure 11 compares the spectral content of the predicted and reference trajectories.
@@
 A total of 1250 distinct airfoil shapes are synthesized by randomly sampling parameters $m\in[0.01,0.09]$, $p\in[0.1,0.7]$, and $t\in[0.1,0.4]$.
+Of these, 1000 samples are used for training and 250 for testing.
 This dataset captures a wide range of geometrical variations and curvature profiles, providing a rigorous benchmark for learning PDE solution operators from geometric input.
-Figure 13: Comparison of predicted signed distance functions (SDFs) for a representative airfoil geometry for Eikonal Equation.
+On this benchmark, the baseline DeepONet attains an average relative $L^{2}$ error of 2.15%, while the proposed FEDONet achieves an improved 1.12% error, demonstrating substantially higher reconstruction accuracy.
+Figure 12: Solving a parametric Eikonal equation (airfoils): Comparison of predicted signed distance functions (SDFs) for a representative airfoil geometry.
 Relative $L^{2}$ error: $2.728\times 10^{-2}$ for DeepONet, $0.912\times 10^{-2}$ for FEDONet.
@@
 This highlights its potential as a general-purpose surrogate for PDEs defined on complex, spatially varying domains.
-3.5 Lorenz-96 Chaotic System
-To assess model performance on nonlinear dynamical systems, we consider the Lorenz-96 (L96) model, a prototypical chaotic system widely used for testing data-driven surrogates in climate modeling and nonlinear forecasting.
-The governing equations are
-$\frac{dx_{i}}{dt}=(x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F,\quad i=1,2,\dots,N,$ (26)
-subject to periodic boundary conditions $x_{i-N}=x_{i}=x_{i+N}$.
-Here, $x_{i}$ denotes the state of the $i$-th variable, $N$ is the system dimension, and $F$ is a constant forcing parameter that determines the level of nonlinearity and chaos.
-For $F>8$, the system exhibits strong chaotic behavior; in this study, we set $F=4.0$ to obtain moderately nonlinear yet stable dynamics.
-The system is integrated using a classical fourth-order Runge–Kutta (RK4) scheme with a fixed time step $\Delta t=0.01$:
-$\mathbf{x}_{n+1}=\mathbf{x}_{n}+\frac{\Delta t}{6}(k_{1}+2k_{2}+2k_{3}+k_{4}),$ (27)
-where $k_{1},k_{2},k_{3},k_{4}$ are the intermediate RK stages computed from the Lorenz–96 derivatives.
-A dataset of 10,00010{,}000 independent trajectories is generated for training and evaluation. Each trajectory is initialized as
-$\mathbf{x}_{0}=F\cdot\mathbf{1}+\epsilon\,\mathcal{N}(0,I),\quad\epsilon=10^{-3},$
-with $N=40$ and integrated for a total physical time of 15 s.
-The initial 10 s are discarded as a transient, and the remaining 5 s (501 time steps) are retained to capture quasi-stationary behavior.
-The resulting dataset has dimensions (10,000,501,40)(10{,}000,501,40), representing 10,000 spatio-temporal trajectories of 40-dimensional state vectors.
-Figure 14: Lorenz-96 - Complete spatio-temporal field reconstruction comparison between DeepONet and FEDONet.
-Figure 14 shows representative test trajectories from the L96 dataset.
-Each panel depicts the initial condition, reference evolution, predictions from the baseline DeepONet and the proposed Fourier-Embedded DeepONet (FEDONet), and their respective absolute-error fields.
-Both models reproduce the characteristic diagonal stripe patterns associated with wave-like phase propagation, maintaining amplitude and wavelength stability over the full 500-step forecast horizon.
-No visible drift or blow-up is observed in either solution.
-Quantitatively, the residual fields indicate comparable performance.
-In the first trajectory, the vanilla DeepONet attains a root-mean-square (RMS) error on the order of $\mathcal{O}(10^{-1})$, while FEDONet slightly reduces this magnitude; in the second trajectory, the ordering reverses locally, but the overall difference remains small.
-On average, FEDONet achieves marginally lower global $\ell_{2}$ errors, though the improvement is less pronounced than in the Poisson or Burgers benchmarks.
-This outcome reflects the inherent structure of the L96 dynamics: the solution manifold is dominated by a slowly varying, low-dimensional spatio-temporal mode, which a shallow fully connected trunk can already approximate effectively.
-Consequently, introducing random Fourier features provides limited additional benefit, consistent with the quantitative results in Table 1, where both models achieve sub-25% relative error.
 3.5 Allen–Cahn Equation
-We evaluate the performance of the proposed Fourier-Embedded DeepONet (FEDONet) in modeling the one-dimensional Allen–Cahn equation nonlinear reaction-diffusion system widely used to describe phase separation processes in multi-component systems.
+We further evaluate the performance of the proposed FEDONet in modeling the one-dimensional Allen–Cahn nonlinear reaction-diffusion system widely used to describe phase separation processes in multi-component systems.
 The governing partial differential equation is given by:
@@
 The goal is to approximate the solution operator $\mathcal{G}\colon s(x)\mapsto u(x,t)$, comparing the performance of vanilla DeepONet and FEDONet.
-Figure 15: Best-performing test case for the Allen–Cahn equation. Relative $\ell_{2}$ error: $4.85\%$ for DeepONet, $2.27\%$ for FEDONet.
-Figure 15 shows qualitative predictions for the best-performing test case.
+Figure 13: Solving a parametric Allen Cahn equation: Comparison of predicted spatio-temporal field for a random initial condition. Relative $L^{2}$ error: $4.85\%$ for DeepONet, $2.27\%$ for FEDONet.
+Both models were evaluated on $128$ randomly selected unseen test samples, yielding average relative $L^{2}$ errors of $10.61\%$ for the DeepONet and $5.50\%$ for the FEDONet.
+Figure 13 shows qualitative predictions for a representative test case.
 FEDONet sharply resolves the interface zones and phase transitions, while DeepONet produces smoother approximations with notable blurring near discontinuities.
-The spectral inductive bias introduced by randomized Fourier features enables FEDONet to retain high-frequency details crucial to phase-field dynamics.
 The pointwise absolute error plot in Figure 14 emphasizes the localized improvements of FEDONet in capturing steep interfaces.
 DeepONet accumulates substantial errors in regions with high spatial gradients, whereas FEDONet maintains predictive accuracy throughout the domain, suggesting better handling of dissipative nonlinearities.
-Figure 16: Allen-Cahn - Energy Spectrum and Pointwise absolute error across space and time for a representative sample.
+Figure 14: Energy spectrum and pointwise absolute error across space and time for a representative sample for Allen-Cahn Equation.
 Figure 14 compares the spatial energy spectra of predicted and true solutions.
 FEDONet accurately tracks the ground truth spectrum across all wavenumbers, preserving the high-frequency tail associated with sharp phase boundaries.
-In contrast, DeepONet suffers from spectral attenuation at $k>20$, revealing an inability to fully capture dissipative and oscillatory dynamics inherent in the Allen–Cahn system.
-These results collectively demonstrate that the incorporation of spectral priors via Fourier embeddings in FEDONet enhances both pointwise accuracy and spectral fidelity.
-The improved resolution of phase interfaces and better retention of high-frequency information position FEDONet as a robust surrogate model for nonlinear, multiscale PDEs such as the Allen–Cahn equation.
+In contrast, DeepONet suffers from spectral attenuation at $k>20$, revealing an inability to fully capture dissipative and oscillatory dynamics inherent in the Allen-Cahn system.
 3.6 Kuramoto–Sivashinsky Equation
-We evaluate the performance of FEDONet on the one-dimensional Kuramoto–Sivashinsky (KS) equation, a canonical model that exhibits spatiotemporal chaos due to the combined effects of nonlinear advection, destabilizing diffusion, and stabilizing fourth-order hyperviscosity.
+We evaluate the performance of FEDONet on the one-dimensional Kuramoto-Sivashinsky (KS) equation, a canonical model that exhibits spatiotemporal chaos due to the combined effects of nonlinear advection, destabilizing diffusion, and stabilizing fourth-order hyperviscosity.
 The governing equation is given by:
@@
 This problem serves as a stringent benchmark due to the sensitivity of chaotic systems to initial conditions and their broadband frequency content.
-Figure 17: Median-performing KS trajectory. Relative $L^{2}$ error: FEDONet = 16.36%, DeepONet = 89.63%.
+Figure 15: Solving a parametric KS equation: Median performing test-sample. Relative $L^{2}$ error: FEDONet = 16.36%, DeepONet = 89.63%.
 Figure 15 shows the reconstructed spatiotemporal field for a median test case.
@@
 Such performance underscores the model’s ability to retain energy in high-frequency modes essential for modeling non-periodic fine-grained behavior that typifies chaotic fields.
-Figure 18: Best-performing KS test sample. FEDONet retains the amplitude and frequency content of the chaotic trajectory, yielding a low relative $L^{2}$ error of $10.77\%$.
-Figure 19: Worst-performing KS test sample. Even in failure cases, FEDONet remains qualitatively reasonable. Relative $L^{2}$ error: FEDONet = $58.02\%$, versus near-complete divergence for DeepONet.
+Figure 16: Solving a parametric KS equation: Best performing test sample.
+FEDONet retains the amplitude and frequency content of the chaotic trajectory, yielding a low relative $L^{2}$ error of $10.77\%$.
+Figure 17: Solving a parametric KS equation: Worst performing test sample.
+Even in failure cases, FEDONet remains qualitatively reasonable.
+Relative $L^{2}$ error: FEDONet = $58.02\%$, versus near-complete divergence for DeepONet.
 Figure 17 displays the worst-case test result.
@@
 This aligns with observations from kernel learning and feature engineering literature, where frequency-based transformations are known to facilitate smoother descent landscapes and gradient propagation.
-Figure 20: KS Training convergence comparison. FEDONet achieves faster and more stable optimization due to implicit spectral regularization.
+Figure 18: KS Training convergence comparison. FEDONet achieves faster and more stable optimization.
 Figure 19: Distribution of relative $L^{2}$ errors across 512 unseen test samples for KS Equation.
@@
 4 Summary
-In this work, we propose the Fourier-Embedded DeepONet (FEDONet) as a principled extension to the classical DeepONet architecture for learning nonlinear operators governed by partial differential equations (PDEs).
+In this work, we propose the Fourier-Embedded DeepONet (FEDONet) as a principled extension to the classical DeepONet architecture for learning nonlinear operators governed by partial differential equations.
 By injecting randomized Fourier features into the trunk input, FEDONet expands the hypothesis space of DeepONets, offering enhanced spectral representation, improved numerical conditioning, and faster convergence during training all without increasing runtime complexity.
 FEDONet consistently outperforms the vanilla DeepONet.
-Notably, it delivers dramatic reductions in relative $\ell_{2}$ error, most prominently in multiscale and high-frequency scenarios such as the Kuramoto-Sivashinsky equation (76% $\rightarrow$ 21%).
-Table 1: Relative $\ell_{2}$ error (%) for each benchmark.
+Notably, it delivers dramatic reductions in relative $L^{2}$ error, most prominently in multiscale and high-frequency scenarios such as the Kuramoto-Sivashinsky equation (76% $\rightarrow$ 21%).
+Table 1: Relative $L^{2}$ error (%) for each benchmark.
 Dataset DeepONet FEDONet
-2-D Poisson (elliptic) 5.80 2.32
-1-D Burgers (conv./diff.) 4.01 2.63
-Lorenz–63 (chaotic ODE) 2.03 0.46
-Lorenz–96 (chaotic ODE) 23.40 22.90
-Allen–Cahn (phase–field) 10.61 5.50
-Kuramoto–Sivashinsky (spatio-chaos) 75.11 20.62
+2-D Poisson (elliptic) 1.41 1.09
+1-D Burgers (conv./diff.) 6.43 4.86
+Lorenz-63 (chaotic ODE) 2.03 0.46
+Allen-Cahn (phase–field) 10.61 5.50
+Kuramoto-Sivashinsky (spatio-chaos) 75.11 20.62
 Eikonal (Hamilton–Jacobi) 2.15 1.12
@@
 Wavelets, Chebyshev polynomials, or orthogonal polynomials tailored to specific PDE families could offer improved localization or boundary adaptivity, especially for systems exhibiting sharp interfaces or multi-scale dynamics.
-Additionally, hybridizing spectral embeddings with structured architectures such as graph-based trunks or encoder–decoder networks may further boost performance on irregular domains or hierarchical solution spaces.
+Additionally, hybridizing spectral embeddings with structured architectures such as graph-based trunks or encoder-decoder networks may further boost performance on irregular domains or hierarchical solution spaces.
 Another avenue involves making the Fourier embedding parameters learnable.
@@
 Data supporting the findings of this study are available from the corresponding author upon reasonable request.
+Declaration of Generative AI use
+During the preparation of this work, the authors used ChatGPT and Grammarly to improve readability and language.
+After using these tools, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
 Appendix A Whitening Effect of Fourier Feature Embeddings
-Let $\phi(\zeta)=\sqrt{2}\,[\sin(2\pi B\zeta),\cos(2\pi B\zeta)]\in\mathbb{R}^{2M}$ be the Fourier feature embedding of an input $\zeta\in\mathbb{R}^{d}$, where each row of the matrix $B\in\mathbb{R}^{M\times d}$ is sampled i.i.d.
+Let $\phi(\zeta)=\sqrt{2}\,[\sin(2\pi Z\zeta),\cos(2\pi Z\zeta)]\in\mathbb{R}^{2M}$ be the Fourier feature embedding of an input $\zeta\in\mathbb{R}^{d}$, where each row of the matrix $Z\in\mathbb{R}^{M\times d}$ is sampled i.i.d.
 from a Gaussian distribution:
-$B_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (32)
+$Z_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (30)
 Assume further that the input $\zeta$ is uniformly distributed over a compact domain, e.g., $\zeta\sim\mathcal{U}([0,1]^{d})$.
@@
 $\mathbb{E}_{\zeta}[\phi(\zeta)\phi(\zeta)^{\top}]\approx I_{2M}.$ (31)
-To see this, consider a single frequency row vector $B_{i}\in\mathbb{R}^{1\times d}$ drawn from $B$. The corresponding $2$-dimensional block of the embedding is:
-$\phi_{i}(\zeta)=\sqrt{2}\begin{bmatrix}\sin(2\pi B_{i}\cdot\zeta)\\ \cos(2\pi B_{i}\cdot\zeta)\end{bmatrix}$. (34)
-We now compute the second-order statistics of this embedding block under the assumption that $B_{i}\cdot\zeta$ is uniformly distributed over $[0,1]$ (a reasonable approximation due to the randomness in $B_{i}$ and the uniformity of $\zeta$).
+To see this, consider a single frequency row vector $Z_{i}\in\mathbb{R}^{1\times d}$ drawn from $Z$. The corresponding $2$-dimensional block of the embedding is:
+$\phi_{i}(\zeta)=\sqrt{2}\begin{bmatrix}\sin(2\pi Z_{i}\cdot\zeta)\\ \cos(2\pi Z_{i}\cdot\zeta)\end{bmatrix}$. (32)
+We now compute the second-order statistics of this embedding block under the assumption that $Z_{i}\cdot\zeta$ is uniformly distributed over $[0,1]$ (a reasonable approximation due to the randomness in $Z_{i}$ and the uniformity of $\zeta$).
 Using standard trigonometric integrals, we obtain:
-$\mathbb{E}_{\zeta}[\sin^{2}(2\pi B_{i}\cdot\zeta)]$ $\approx\frac{1}{2},$ (35)
-$\mathbb{E}_{\zeta}[\cos^{2}(2\pi B_{i}\cdot\zeta)]$ $\approx\frac{1}{2},$ (36)
-$\mathbb{E}_{\zeta}[\sin(2\pi B_{i}\cdot\zeta)\cos(2\pi B_{i}\cdot\zeta)]$ $\approx 0.$ (37)
+$\mathbb{E}_{\zeta}[\sin^{2}(2\pi Z_{i}\cdot\zeta)]$ $\approx\frac{1}{2},$ (33)
+$\mathbb{E}_{\zeta}[\cos^{2}(2\pi Z_{i}\cdot\zeta)]$ $\approx\frac{1}{2},$ (34)
+$\mathbb{E}_{\zeta}[\sin(2\pi Z_{i}\cdot\zeta)\cos(2\pi Z_{i}\cdot\zeta)]$ $\approx 0.$ (35)
 Therefore, the unscaled covariance matrix for this block is approximately:
-$\mathbb{E}_{\zeta}\left[\begin{bmatrix}\sin(2\pi B_{i}\cdot\zeta)\\ \cos(2\pi B_{i}\cdot\zeta)\end{bmatrix}\begin{bmatrix}\sin(2\pi B_{i}\cdot\zeta)&\cos(2\pi B_{i}\cdot\zeta)\end{bmatrix}\right]\approx\frac{1}{2}I_{2}.$ (38)
+$\mathbb{E}_{\zeta}\left[\begin{bmatrix}\sin(2\pi Z_{i}\cdot\zeta)\\ \cos(2\pi Z_{i}\cdot\zeta)\end{bmatrix}\begin{bmatrix}\sin(2\pi Z_{i}\cdot\zeta)&\cos(2\pi Z_{i}\cdot\zeta)\end{bmatrix}\right]\approx\frac{1}{2}I_{2}.$ (36)
 Scaling by the $\sqrt{2}$ factor yields:
 $\mathbb{E}_{\zeta}[\phi_{i}(\zeta)\phi_{i}(\zeta)^{\top}]\approx I_{2}.$ (37)
-Since the rows of $B$ are sampled independently and the embeddings for each $B_{i}$ are uncorrelated, the full embedding vector $\phi(\zeta)\in\mathbb{R}^{2M}$ satisfies:
+Since the rows of $Z$ are sampled independently and the embeddings for each $Z_{i}$ are uncorrelated, the full embedding vector $\phi(\zeta)\in\mathbb{R}^{2M}$ satisfies:
 $\mathbb{E}_{\zeta}[\phi(\zeta)\phi(\zeta)^{\top}]\approx I_{2M}.$ (38)
@@
 Let $\mathcal{H}_{\text{vanilla}}$ be the hypothesis space of functions representable by a feedforward neural network with input $\zeta\in\mathbb{R}^{d}$, and let $\mathcal{H}_{\phi}$ be the hypothesis space of functions representable by a similar network where the input is first mapped via the Fourier feature embedding:
-$\phi(\zeta)=\left[\sin(2\pi B\zeta),\cos(2\pi B\zeta)\right],\quad B\in\mathbb{R}^{M\times d},\;B_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (41)
+$\phi(\zeta)=\left[\sin(2\pi Z\zeta),\cos(2\pi Z\zeta)\right],\quad Z\in\mathbb{R}^{M\times d},\;Z_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (39)
 Then, under mild regularity assumptions (e.g., compact input domain $K\subset\mathbb{R}^{d}$ and continuous activation functions), we have:
@@
 Due to the spectral bias of neural networks [51], this high-frequency function is difficult to approximate using shallow networks on raw input.
-However, if the frequency matrix $B$ used in the embedding $\phi(\zeta)$ includes rows close to $100$, then:
-$\sin(2\pi B_{j}\zeta)\approx\sin(100\pi\zeta),$
+However, if the frequency matrix $Z$ used in the embedding $\phi(\zeta)$ includes rows close to $100$, then:
+$\sin(2\pi Z_{j}\zeta)\approx\sin(100\pi\zeta),$
 making $f$ easily representable in $\mathcal{H}_{\phi}$.
@@
   3. 3.
-    The operator learning framework parallels well-established numerical discretization techniques, such as Petrov–Galerkin and spectral methods, but operates in a fully data-driven regime.
+    The operator learning framework parallels well-established numerical discretization techniques, such as Petrov-Galerkin and spectral methods, but operates in a fully data-driven regime.
 In summary, Fourier Embeddings serve as a principled mechanism for spectral lifting, enabling DeepONet to perform a learned Galerkin-style decomposition of operators.

### v4

@@
 Despite their promising capabilities, the standard implementation of DeepONets, which typically employs fully connected linear layers in the trunk network, can encounter limitations in capturing complex spatial structures inherent to various PDEs.
-To address this limitation, we introduce Fourier-Embedded trunk networks within the DeepONet architecture, leveraging random fourier feature mappings to enrich spatial representation capabilities.
-Our proposed Fourier-Embedded DeepONet, FEDONet demonstrates superior performance compared to the traditional DeepONet across a comprehensive suite of PDE-driven datasets, including the two-dimensional Poisson, Burgers’, Lorenz-63, Eikonal, Allen-Cahn, and the Kuramoto-Sivashinsky equation.
+To address this limitation, we use Fourier-Embedded trunk networks within the DeepONet architecture, leveraging random Fourier features to enrich spatial representation capabilities.
+The Fourier-Embedded DeepONet (FEDONet) demonstrates superior performance compared to the traditional DeepONet across a comprehensive suite of PDE-driven datasets, including the Burgers’, 2D Poisson, Eikonal, Allen-Cahn, and the Kuramoto-Sivashinsky equation.
+To systematically evaluate the effectiveness of the architectures, we perform comparisons across multiple training dataset sizes and input noise levels.
 FEDONet delivers consistently superior reconstruction accuracy across all benchmark PDEs, with particularly large relative $L^{2}$ error reductions observed in chaotic and stiff systems.
-This study highlights the effectiveness of Fourier embeddings in enhancing neural operator learning, offering a robust and broadly applicable methodology for PDE surrogate modeling.
+This work demonstrates the effectiveness of Fourier embeddings in enhancing neural operator learning, offering a robust and broadly applicable methodology for PDE surrogate modeling.
 keywords:
@@
 […]
-organization=Mechanical and Aerospace Engineering, addressline=1512 Middle Drive, University of Tennessee, city=Knoxville, postcode=37996-2210, state=TN, country=USA
+[aff1]organization=Mechanical and Aerospace Engineering, addressline=1512 Middle Drive, University of Tennessee, city=Knoxville, postcode=37996-2210, state=TN, country=USA
 1 Introduction
 Partial differential equations underpin models ranging from turbulent‐flow dynamics and heat diffusion to electromagnetics and biological transport.
-Classical numerical solvers such as finite‐difference, finite‐element and spectral methods provide high‐fidelity solutions but incur a prohibitive computational burden when faced with high‐dimensional, multiscale or stiff systems [1, 2, 3].
-Efforts to construct efficient surrogate models have a long history, spanning the universal approximation theorems for neural networks[4, 5, 6, 7], reduced‐order techniques based on proper‐orthogonal decomposition [8, 9], kernel Gaussian process surrogates [10], and mesh‐free radial‐basis schemes [11, 12].
-Early data‐driven approaches also sought to learn governing equations directly from trajectory data using deep neural networks [13] However, extending these approaches to learn function‐to‐function mappings capable of generalizing across discretizations, boundary conditions, and input configurations has remained a long standing challenge.
+Classical numerical solvers such as finite‐difference, finite‐element, and spectral methods provide high‐fidelity solutions but incur a prohibitive computational burden when faced with high‐dimensional, multiscale or stiff systems [1, 2, 3].
+Efforts to construct efficient surrogate models have a long history, spanning the universal approximation theorems for neural networks [4, 5, 6, 7], reduced‐order techniques based on proper‐orthogonal decomposition [8, 9], kernel-based Gaussian process surrogates [10], and mesh‐free radial‐basis schemes [11, 12].
+Early data‐driven approaches also sought to learn governing equations directly from trajectory data using deep neural networks [13].
+However, extending these approaches to learn function‐to‐function mappings capable of generalizing across discretizations, boundary conditions, and input configurations has remained a long-standing challenge.
+From a numerical analysis perspective, this difficulty can often be interpreted as a mismatch between the spectral content of PDE solutions and the coordinate representations supplied to standard neural architectures.
+In such cases, the network must learn highly oscillatory or multiscale structures solely through nonlinear transformations, which can lead to inefficient optimization and poor sample efficiency.
 The recent emergence of operator learning addresses this long-standing challenge by training neural networks that act directly on infinite-dimensional function spaces, enabling generalization across discretizations and grid resolutions.
@@
 Subsequent developments have further advanced this paradigm through multiwavelet representations [20, 21], convolutional kernels [22], multigrid tensorizations [23, 24], and geometry-aware spectral deformations [25], collectively enhancing scalability, resolution fidelity, and geometric adaptability.
-Beyond architectural and kernel innovations, recent research has expanded operator learning along multiple frontiers.
+Despite these advances, a key challenge remains the development of operator learning architectures that balance spectral accuracy, computational efficiency, and architectural simplicity, particularly for integration into large-scale scientific workflows such as turbulence modeling and data assimilation in fluid mechanics.
+Recent research has expanded operator learning along multiple frontiers.
 Efforts addressing solution irregularities have focused on handling discontinuities [26], incorporating derivative supervision for enhanced gradient fidelity [27], and enabling learning on irregular meshes through graph-based trunk networks [28].
@@
 Parallel efforts have explored latent and invertible operator formulations to achieve data-efficient and bidirectional inference [38, 39, 40, 41].
-Transformer based attention mechanisms have also been adapted for operator learning [42, 43, 44, 45, 46] and for robust geometry generalization [47].
+Transformer-based attention mechanisms have also been adapted for operator learning [42, 43, 44, 45, 46] and for robust geometry generalization [47].
 Together, these advances have enabled high-fidelity, climate-scale surrogates such as FourCastNet [48] and its amortized Fourier Neural Operator successor [49], underscoring the maturity and scalability of the operator-learning paradigm.
-Despite the recent operator-learning architectures, the trunk network of conventional DeepONets remains a shallow multilayer perceptron.
+Recent efforts have also begun to explore the application of data-driven operator learning and spectral modeling approaches to more complex fluid dynamics systems, highlighting the importance of robustness, spectral fidelity, and generalization in realistic regimes [50, 51].
+Despite recent operator-learning architectures, the trunk network of conventional DeepONets remains a shallow multilayer perceptron.
 This design often struggles to represent the oscillatory and multi-scale structures that characterize nonlinear PDE solutions, resulting in reduced sample efficiency and reconstruction accuracy.
-While recent alternatives such as graph-based [28] or wavelet-enhanced U-Net trunks [30] offer improved expressivity, they introduce significant architectural complexity and domain-specific preprocessing requirements.
-In contrast, random Fourier features have proven remarkably effective at encoding high-frequency information in standard deep networks [50], yet their potential within operator-learning frameworks remains largely unexplored, where the trunk is typically implemented as a simple fully connected MLP.
-To address this gap, we propose the Fourier-Embedded Deep Operator Network (FEDONet), which augments the DeepONet trunk by replacing raw spatial coordinates with a compact random Fourier-feature embedding [50].
-This embedding implicitly introduces a spectrum of sinusoidal basis functions, enriching the trunk’s representational capacity with negligible additional parameters.
-FEDONet retains the original branch-trunk separation, integrates seamlessly with existing DeepONet implementations, and incurs no runtime overhead beyond a single matrix multiplication.
-We formally introduce the Fourier embedding of trunk coordinates and analyze its theoretical implications through the lens of the operator neural tangent kernel, demonstrating that the embedding strictly enlarges the approximation class of standard DeepONets.
-A comprehensive benchmark is carried spanning seven canonical PDE families - steady Poisson, viscous Burgers, Allen-Cahn, Kuramoto-Sivashinsky, Eikonal, and Lorenz-63.
-By embedding a principled spectral bias into the trunk network, FEDONet bridges the performance gap to frequency-domain operators, while preserving the architectural simplicity, flexibility, and locality benefits of the original DeepONet formulation.
-Section  2 presents the formulation of the Fourier-Embedded Deep Operator Network (FEDONet), detailing the embedding construction, its theoretical connection to the operator neural tangent kernel, and its implications for approximation capacity.
-Section  3 describes the benchmark setup and implementation details across the seven PDE families considered.
-Sections  4 and  5 summarize the main findings and outlines directions for future research in spectral operator learning and physics-constrained network design.
+While recent alternatives such as graph-based [28] or wavelet-enhanced U-Net-based trunks [30] offer improved expressivity, they introduce significant architectural complexity and domain-specific preprocessing requirements.
+In contrast, random Fourier features have proven remarkably effective at encoding high-frequency information in standard deep networks [52], yet their potential within operator-learning frameworks remains largely unexplored, particularly in settings where the trunk is typically implemented as a simple fully connected MLP.
+From this viewpoint, enriching the coordinate representation with sinusoidal embeddings can be interpreted as a form of spectral lifting that aligns the input representation with the frequency content of typical PDE solutions.
+Recent studies have also explored incorporating Fourier-based representations within DeepONet frameworks for surrogate modeling in specific application domains [53].
+In contrast to such approaches that rely on predefined Fourier bases tailored to particular physical settings, the present work introduces randomized Fourier feature embeddings applied directly to the trunk coordinates.
+This strategy acts as a lightweight and general-purpose coordinate transformation that enriches the spectral representation capacity of the trunk network while preserving the standard DeepONet formulation.
+The Fourier-Embedded DeepONet, FEDONet retains the original branch-trunk separation, integrates seamlessly with existing DeepONet implementations, and incurs no runtime overhead beyond a single matrix multiplication.
+We formally introduce the Fourier embedding of trunk coordinates and analyze its theoretical implications through the lens of the operator neural tangent kernel, highlighting its role as a spectral preconditioning mechanism that improves optimization behavior and training stability.
+This analysis demonstrates that the embedding strictly enlarges the approximation class of standard DeepONets.
+We conduct a comprehensive benchmark spanning steady 2D Poisson, 1D Viscous Burgers, 1D Allen-Cahn, 1D Kuramoto-Sivashinsky, and 2D Eikonal equation.
+In addition to accuracy comparisons, we perform systematic analyses examining the dependence of model performance on training dataset size and input noise levels, providing further insight into the robustness and data efficiency of the architectures.
+By embedding a principled spectral bias into the trunk network, FEDONet bridges the performance gap to frequency-domain operators while preserving the architectural simplicity, flexibility, and locality benefits of the original DeepONet formulation.
+The primary contributions of this work are summarized as follows.
+First, we introduce a Fourier-Embedded Deep Operator Network (FEDONet) that enriches trunk coordinate representations through randomized Fourier feature embeddings, providing a lightweight spectral lifting mechanism for DeepONet architectures.
+Second, we provide theoretical insight into the effect of this embedding through the operator neural tangent kernel, showing that the resulting architecture expands the effective hypothesis space of standard DeepONets while improving optimization conditioning.
+Third, we perform a systematic empirical evaluation across multiple canonical PDE benchmarks and analyze robustness with respect to dataset size and input noise levels.
+Together, these results demonstrate that spectral coordinate embeddings provide a simple yet effective mechanism for improving operator learning performance without increasing architectural complexity.
 2 Methodology
 We aim to develop a data-driven framework that approximates nonlinear operators mapping between infinite-dimensional function spaces.
-Let $\Omega\subset\mathbb{R}^{D}$ be a bounded domain, and define the input and output function spaces:
+Let $\Omega\subset\mathbb{R}^{D}$ be a bounded domain, and define the input and output function spaces,
 $\mathcal{U}$ $=\left\{u:\mathcal{X}\rightarrow\mathbb{R}^{d_{u}}\right\},\quad\mathcal{X}\subseteq\mathbb{R}^{d_{x}},$ (1)
@@
 Our goal is to learn an approximation $\mathcal{G}_{\theta}:\mathcal{U}\rightarrow\mathcal{S}$ to an unknown operator $\mathcal{G}$, where $\theta\in\Theta$ are trainable parameters.
-In practice, we are given a dataset $\mathcal{D}=\{(u^{i},s^{i})\}_{i=1}^{N}$, comprising function pairs sampled from $\mathcal{G}$.
+In practice, we are given a dataset $\mathcal{D}=\{(u^{i},s^{i})\}_{i=1}^{N}$ comprising function pairs sampled from $\mathcal{G}$.
 DeepONet is a neural operator architecture well-suited to this task.
-It decomposes the learning problem into two components: Branch network $B_{\theta}$ that encodes discrete samples of $u$ at sensor points $\{x_{1},\dots,x_{m}\}$, Trunk network $T_{\theta}$ that takes in spatial (or spatiotemporal) coordinates $\zeta=(x,y,z,t)$.
-The predicted value at location $\zeta$ is computed via an inner product:
+It decomposes the learning problem into two components,
+  • Branch network $B_{\theta}$ that encodes discrete samples of the input function $u$ evaluated at sensor locations $\{x_{1},\dots,x_{m}\}$.
+  • Trunk network $T_{\theta}$ that receives spatial or spatiotemporal coordinates $\zeta\in\mathcal{Y}$ (e.g., $\zeta=(x,y,z,t)$).
+The predicted value of the operator at location $\zeta$ is computed through an inner product between the branch and trunk outputs,
 $\mathcal{G}_{\theta}(u)(\zeta)=B_{\theta}(u)\cdot T_{\theta}(\zeta).$ (3)
 2.1 Fourier Embeddings as Spectral Preconditioners
-Standard MLPs exhibit a well-known spectral bias - they learn low-frequency components faster than high-frequency ones [51].
-This bottleneck is particularly harmful in operator learning problems involving sharp gradients, multiscale dynamics, or oscillatory fields.
-To address this, we propose injecting fixed Fourier Embeddings into the trunk network input illustrated in Fig.
+Standard multilayer perceptrons (MLPs) exhibit a well-known spectral bias, whereby low-frequency components are learned significantly faster than high-frequency components [54].
+This phenomenon becomes particularly problematic in operator learning tasks involving sharp gradients, multiscale dynamics, or oscillatory solution fields.
+From a numerical perspective, this limitation can be viewed as a conditioning issue arising from an inadequate coordinate representation.
+When raw spatial coordinates are used as inputs, the network must internally construct oscillatory features through deep nonlinear transformations, which can slow optimization and reduce sample efficiency.
+To address this limitation, we use fixed Fourier Embeddings applied to the trunk network input, illustrated in Fig.
 1.
@@
 $\phi(\zeta)=\left[\sin(2\pi Z\zeta),\cos(2\pi Z\zeta)\right],\quad Z_{ij}\sim\mathcal{N}(0,\sigma^{2}),$ (4)
-where $Z\in\mathbb{R}^{M\times d}$ is a Gaussian frequency matrix. Instead of feeding raw coordinates to the trunk, we input $\phi(\zeta)$ - enriching its spectral content.
-The FEDONet prediction then becomes:
-$\mathcal{G}_{\theta}(u)(\zeta)=B_{\theta}(u)\cdot T_{\theta}(\phi(\zeta))=\sum_{k=1}^{p}b_{k}(u)\cdot t_{k}(\phi(\zeta)).$ (5)
-Figure 1: FEDONet: Fourier Embedded Deep operator Network
-This transformation can be interpreted as a randomized kernel approximation:
+where $Z\in\mathbb{R}^{M\times d}$ is a Gaussian frequency matrix.
+Instead of feeding raw coordinates to the trunk network, we provide the transformed features $\phi(\zeta)$, thereby enriching the spectral representation of the coordinate inputs.
+Unlike Fourier-basis DeepONet [53] approaches that construct the input function space using predefined Fourier modes, the present formulation does not impose any explicit spectral structure on the input functions themselves.
+Instead, Fourier embeddings act purely as a coordinate transformation that enhances the representational capacity of the trunk network while preserving the standard DeepONet architecture.
+The FEDONet prediction therefore becomes
+$\mathcal{G}_{\theta}(u)(\zeta)=B_{\theta}(u)\cdot T_{\theta}(\phi(\zeta))=\sum_{k=1}^{p}b_{k}(u)\,t_{k}(\phi(\zeta)).$ (5)
+Figure 1: FEDONet: Fourier-Embedded Deep Operator Network.
+This transformation can be interpreted as a randomized kernel approximation
 $k(\zeta,\zeta^{\prime})\approx\phi(\zeta)^{\top}\phi(\zeta^{\prime}),$ (6)
-effectively lifting $\zeta$ into a high-dimensional basis of sinusoidal functions.
-By Bochner’s theorem [52], such embeddings approximate shift-invariant kernels - implicitly kernelizing the trunk network.
-This expands its representational capacity, enabling better learning of high-frequency targets.
-Beyond expressivity, Fourier embeddings improve the conditioning of the learning process. Under mild assumptions, the embedded features exhibit approximate whitening ( see A ) :
+which lifts the coordinate $\zeta$ into a high-dimensional basis of sinusoidal functions.
+By Bochner’s theorem [55], such embeddings approximate shift-invariant kernels and therefore implicitly kernelize the trunk network.
+This spectral lifting enables the trunk network to represent oscillatory and multiscale spatial patterns using shallow architectures, improving the learning efficiency for PDE operators with rich frequency content.
+Beyond representational expressivity, Fourier embeddings also improve the conditioning of the learning problem. Under mild assumptions, the embedded features exhibit approximate whitening (see A):
 $\mathbb{E}_{\zeta}[\phi(\zeta)\phi(\zeta)^{\top}]\approx I,$ (7)
-reducing feature correlations and leading to a more isotropic Neural Tangent Kernel (NTK) spectrum.
-This in turn improves optimization dynamics, stabilizes gradient flow, and accelerates convergence during training.
-Let $\mathcal{H}_{\phi}$ denote the function class learnable with Fourier features, and $\mathcal{H}_{\text{vanilla}}$ the class learnable with raw coordinates. Then:
-$\mathcal{H}_{\phi}\supset\mathcal{H}_{\text{vanilla}},$ (8)
-with the inclusion being strict in problems with rich frequency content (see B). Fourier embeddings thus act as spectral preconditioners that align the input geometry with the task’s spectral demands.
+which reduces feature correlations and produces a more isotropic Neural Tangent Kernel (NTK) spectrum.
+This improved conditioning leads to more stable gradient propagation and faster convergence during training, effectively acting as a spectral preconditioning mechanism for operator learning.
 2.2 Training Objective
@@
 The objective is to approximate a target operator $G$ by minimizing the prediction error over a dataset of paired input-output functions.
-Each dataset element consists of a function $u^{(i)}\in\mathcal{U}$ and corresponding evaluations of the output $s^{(i)}_{j}=G(u^{(i)})(\zeta^{(i)}_{j})$ at query locations $\zeta^{(i)}_{j}\in\mathcal{Y}$.
-Let the dataset be denoted by
-$\mathcal{D}=\left\{\left(u^{(i)},\left\{\zeta^{(i)}_{j},s^{(i)}_{j}\right\}_{j=1}^{Q}\right)\right\}_{i=1}^{N},$
-where $N$ is the number of input functions and $Q$ is the number of query points per function.
-The model prediction at location $\zeta$ is given by:
-$G_{\theta}(u)(\zeta)=\sum_{k=1}^{p}b_{k}(u)\cdot t_{k}(\phi(\zeta)),$
-where $b_{k}(u)$ and $t_{k}(\phi(\zeta))$ denote the outputs of the branch and Fourier-embedded trunk networks, respectively.
-The training objective is to minimize the empirical risk using the mean squared error (MSE) loss:
+Each dataset element consists of an input function $u^{(i)}\in\mathcal{U}$ and corresponding evaluations of the output $s^{(i)}_{j}=G(u^{(i)})(\zeta^{(i)}_{j})$ at query locations $\zeta^{(i)}_{j}\in\mathcal{Y}$.
+The dataset can therefore be written as
+$\mathcal{D}=\left\{\left(u^{(i)},\{\zeta^{(i)}_{j},s^{(i)}_{j}\}_{j=1}^{Q}\right)\right\}_{i=1}^{N}.$
+The model prediction at location $\zeta$ is
+$G_{\theta}(u)(\zeta)=\sum_{k=1}^{p}b_{k}(u)\,t_{k}(\phi(\zeta)).$
+The training objective minimizes the empirical risk using the mean squared error loss:
 $\mathcal{L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{Q}\sum_{j=1}^{Q}\left\|G_{\theta}(u^{(i)})(\zeta^{(i)}_{j})-s^{(i)}_{j}\right\|^{2}.$
-Model parameters are optimized using mini-batch stochastic gradient descent with the Adam optimizer.
-Learning rate schedules are selected based on the problem characteristics to ensure stable and efficient convergence.
-Each mini-batch consists of randomly sampled function-query pairs $(u,\zeta)$, enabling scalable and efficient training across diverse domains.
-Notably, the Fourier embedding layer $\phi(\zeta)$ is fixed and non-trainable throughout training.
-This design choice ensures low overhead while enhancing model expressivity through spectral enrichment.
-Nonetheless, the architecture can be extended to incorporate learnable embeddings $\phi_{\theta}(\zeta)$, enabling adaptive spectral tuning tailored to problem-specific frequency content.
-2.3 Evaluation and Generalization
-After training, the FEDONet model is evaluated on unseen input functions $u_{\text{test}}$. To assess model performance, we report both quantitative and qualitative metrics:
+Model parameters are optimized using mini-batch optimization using Adam.
+Each mini-batch consists of randomly sampled function-coordinate pairs $(u,\zeta)$, which enables scalable training and improves generalization across diverse operator inputs.
+The Fourier embedding layer $\phi(\zeta)$ remains fixed and non-trainable throughout training, ensuring minimal computational overhead while enriching the coordinate representation.
+Although the present work focuses on fixed random embeddings, the framework can naturally be extended to learnable embeddings $\phi_{\theta}(\zeta)$ that adapt the spectral representation to problem-specific frequency content.
+2.3 Evaluation
+After training, DeepONet and FEDONet are evaluated on unseen input functions $u_{\text{test}}$ sampled from the same distribution as the training set.
+The evaluation is designed to probe not only pointwise accuracy but also the model’s ability to capture multiscale structures, remain robust under varying noise levels, and generalize effectively in limited-data regimes.
+To assess model performance, we report both quantitative and qualitative metrics:
   • Relative $L^{2}$ Error: Defined as
-    $\varepsilon_{L^{2}}=\frac{\|G_{\theta}(u)-G(u)\|_{2}}{\|G(u)\|_{2}},$
-    where the norm is computed over a dense evaluation grid. This metric quantifies the normalized discrepancy between predicted and reference output fields.
-  • Spectral Fidelity: We perform Fourier analysis of the predicted and true fields to evaluate preservation of high-frequency content.
-Specifically, we compare the angle-integrated power spectra $E(k)$ across wavenumbers $k$, assessing the model’s ability to capture multiscale dynamics and resolve fine-scale features.
-The ability to evaluate the learned operator at arbitrary coordinates makes FEDONet particularly suitable for scientific computing tasks requiring continuous surrogate modeling.
-In all benchmark scenarios, FEDONet demonstrates strong generalization, enhanced spectral accuracy, and robust performance across a wide range of PDE systems with varying stiffness, nonlinearity, and complexity.
+    $\varepsilon_{L^{2}}=\frac{\|G_{\theta}(u)-G(u)\|_{2}}{\|G(u)\|_{2}}.$
+    where the norm is computed over a dense evaluation grid.
+This metric quantifies the normalized discrepancy between predicted and reference output fields and serves as the primary measure of reconstruction accuracy.
+  • Spectral Fidelity: To evaluate the preservation of multiscale structures, we analyze the Fourier energy spectrum of predicted and reference fields.
+Specifically, we compare the angle-integrated power spectra $E(k)$ across wavenumbers $k$.
+This diagnostic provides a stringent test of the model’s ability to recover high-frequency content, which is often underrepresented in operator learning frameworks.
+Particular emphasis is placed on the decay rate and tail behavior of $E(k)$, which indicate how well fine-scale dynamics are captured.
+  • Robustness to Input Noise: We evaluate model performance under varying levels of additive noise in the input functions,
+    $u_{\text{test}}^{\sigma}=u_{\text{test}}+\sigma\,\eta,$
+    where $\eta$ is Gaussian noise.
+Here, $\sigma$ is expressed as a fraction of the standard deviation of the input signal, and is reported in percentage terms (e.g., $\sigma=0.1$ corresponds to $10\%$ noise).
+The degradation of $\varepsilon_{L^{2}}$ with increasing noise levels provides insight into the stability of the learned operator.
+This analysis is critical for practical scientific applications where measurements are inherently noisy.
+  • Sample Efficiency: To assess how efficiently the model learns the operator, we train FEDONet and baseline DeepONet models using varying numbers of training samples.
+The resulting error scaling with respect to dataset size reveals the data efficiency of the architecture and highlights the impact of Fourier embeddings in low-data regimes.
 3 Results
+3.1 Burgers’ Equation
+The viscous Burgers equation is a canonical 1D nonlinear PDE. It has been studied extensively for DeepONets in prior literature [56]. We consider the governing equation,
+$\frac{\partial u}{\partial t}(x,t)+u\frac{\partial u}{\partial x}(x,t)=\nu\frac{\partial^{2}u}{\partial x^{2}}(x,t),\>\>\>\>\>\forall(x,t)\in[0,1]\times[0,1]$ (8)
+where $x$ and $t$ denote the spatio-temporal coordinates and $\nu$ is the kinematic viscosity. We set $\nu=0.01$ along with initial and periodic boundary conditions as,
+$\begin{split}\displaystyle u(x,0)=&\>s(x),\>\>\>\>\>\>\>\>\>\>\>\forall\>x\in[0,1]\\ u(0,t)=&\>u(1,t),\>\>\>\>\>\>\>\>\forall\>t\in[0,1]\\ \frac{\partial u}{\partial x}(0,t)=&\>\frac{\partial u}{\partial x}(1,t)\>\>\>\>\>\>\forall\>t\in[0,1]\end{split}$ (9)
+where the initial condition $s(x)$ is sampled from a Gaussian Random Field, $s(x)\sim\mathcal{N}\left(0,25^{2}(-\Delta+5^{2}I)^{-4}\right)$, satisfying periodicity.
+The objective is to learn the nonlinear operator $\mathcal{G}$ mapping $s(x)$ to the full spatio-temporal solution $u(x,t)$ using both vanilla DeepONet and the FEDONet.
+Both models are evaluated on $250$ unseen, noise-free test samples, yielding average relative $L^{2}$ errors of $6.43\%$ for DeepONet and $4.86\%$ for FEDONet.
+Figure 2 shows predictions for a test input corrupted with $10\%$ additive noise.
+DeepONet captures the coarse solution structure but fails to resolve sharp gradients, exhibiting pronounced smoothing near shock regions.
+In contrast, FEDONet yields markedly sharper reconstructions, accurately capturing shock fronts and preserving fine-scale dynamics.
+Figure 2: Parametric Burgers’ Equation: Comparison of predicted spatio-temporal solution fields for a test input with additive noise.
+The corresponding relative $L^{2}$ errors are $6.5\%$ for DeepONet and $5.0\%$ for FEDONet.
+The energy spectra of the predicted and ground truth fields, shown in Figure 3, further emphasize this advantage.
+FEDONet accurately captures the spectral decay across a wide range of wavenumbers, aligning with the physical expectations of dissipative transport phenomena.
+DeepONet, however, underestimates energy at higher modes, reflecting its inability to resolve fine features and spectral richness.
+Figure 3: Spectral and Error analysis for Burgers’ equation: Comparison of the energy spectrum and pointwise absolute error fields for a representative noise-free test sample.
+Figure 4 illustrates the training loss convergence for DeepONet and FEDONet across a range of batch sizes.
+Several consistent trends are observed.
+First, FEDONet exhibits significantly faster initial convergence, rapidly reducing the loss within the early training epochs compared to DeepONet.
+Second, across all batch sizes, FEDONet consistently attains lower final loss values, indicating a more accurate approximation of the underlying operator.
+Notably, the performance gap between the two models persists regardless of batch size, demonstrating that the advantage of FEDONet is not merely an artifact of a particular training configuration.
+As the batch size increases, both models exhibit smoother training curves due to reduced stochasticity; however, DeepONet remains systematically biased toward higher loss values and slower convergence.
+Figure 4: Training convergence for Burgers’ equation: Comparison of training loss evolution for DeepONet and FEDONet across different batch sizes (128, 256, 512, 1024).
+Figure 5 examines the impact of the number of Fourier features on FEDONet performance.
+As the number of features increases, the error decreases rapidly initially, indicating improved capacity to capture high-frequency components.
+Beyond a certain threshold, the performance saturates and exhibits mild fluctuations, suggesting a balance between representational richness and over-parameterization.
+Figure 5: Effect of Fourier Feature Resolution: Mean Relative $L^{2}$ error for FEDONet as a function of the number of Fourier features.
 3.2 2D Poisson Equation
-We constructed a synthetic dataset comprising numerical solutions to the canonical two-dimensional Poisson equation. The governing equation is
+The 2D Poisson equation serves as a canonical model for elliptic partial differential equations arising in a wide range of physical systems, including electrostatics, heat conduction, and potential flow.
+We consider the governing equation,
 $\nabla^{2}u(x,y)=f(x,y),\quad(x,y)\in[0,1]^{2},$ (10)
@@
 where $u(x,y)$ denotes the potential field and $f(x,y)$ represents a prescribed source distribution.
-A total of N=10, [^0] N=10{,}000 distinct source fields were synthesized using Gaussian random fields (GRFs) with spectral parameters $\alpha=3$ and $\tau=3$,
+A total of N=10, [^0] N=10{,}000 distinct source fields were synthesized using Gaussian random fields (GRFs) with smoothness parameter $\alpha=3$ and length scale $\tau=3$.
 $f\sim\text{GRF}(\alpha=3,\tau=3),$ (12)
@@
 For each realization of $f$, the corresponding solution $u$ was obtained by numerically solving the Poisson equation using a standard five-point finite-difference discretization of the Laplacian operator.
-The spatial domain $\Omega=[0,1]^{2}$ was thus transformed into a structured algebraic system,
+Discretization over $\Omega=[0,1]^{2}$ yields the structured algebraic system,
 $A\mathbf{u}=\mathbf{b},$ (13)
@@
 The coefficient matrix $A$ was defined using a conventional five-point stencil with Dirichlet conditions enforced by setting boundary rows to the identity and zeroing the corresponding entries in $\mathbf{b}$.
-This process yielded 10,000 pairs of input-output fields suitable for training, validation, and testing of neural operator architectures.
-Figure 2: Solving a 2D Poisson equation: Comparison of predicted and reference solution fields of DeepONet and FEDONet
-Figure 2 shows a representative test case from the Poisson dataset.
-The top row depicts the input forcing field, exact solution, prediction from the baseline DeepONet, and the corresponding pointwise error, while the bottom row presents results from the proposed Fourier-Embedded DeepONet (FEDONet).
-Both models were evaluated on $1000$ randomly selected unseen test samples, yielding average relative $L^{2}$ errors of $1.41\%$ for the DeepONet and $1.09\%$ for the FEDONet.
-FEDONet reproduces the reference field with sharper gradients and more accurate spatial localization of extrema.
-The residual field of the vanilla DeepONet displays fine-scale artifacts, while FEDONet yields a smoother, spatially coherent error distribution, corresponding to an approximate fivefold improvement in reconstruction accuracy.
-Figure 3: Energy spectrum comparison between the ground truth, vanilla and Fourier DeepOnet reconstructions for the 2D Poisson Dataset.
-Spectral accuracy was assessed using the angle-integrated energy spectrum $E_{k}$ shown in Fig.
-3.
+Figure 6: Solving a 2D Poisson equation: Comparison of predicted and reference solution fields of DeepONet and FEDONet for a test input with $10\%$ additive noise.
+The corresponding relative $L^{2}$ errors are $2.85\%$ for DeepONet and $1.73\%$ for FEDONet.
+Figure 6 presents a representative test case from the Poisson dataset.
+The top row shows the input noisy forcing field, the ground truth solution, the DeepONet prediction, and the corresponding pointwise error, while the bottom row displays the same quantities for FEDONet.
+Both models were evaluated on $1000$ randomly selected unseen test samples, yielding average relative $L^{2}$ errors of $1.41\%$ for DeepONet and $1.09\%$ for FEDONet under noise-free input conditions.
+While both models achieve high overall accuracy, FEDONet consistently produces sharper gradients and more precise spatial localization of extrema.
+In contrast, the error field of DeepONet exhibits fine-scale oscillatory artifacts, indicating limitations in resolving localized structures.
+FEDONet, however, yields a smoother and more spatially coherent error distribution, reflecting improved reconstruction quality.
+Figure 7: Spectral analysis for 2D Poisson equation: Energy spectra of the ground truth solution and the corresponding reconstructions obtained using DeepONet and FEDONet for a representative noise-free test sample.
+Spectral accuracy was assessed using the angle-integrated energy spectrum $E_{k}$ shown in Figure 7.
 Both models reproduce the reference spectrum at large scales ($k\lesssim 50$), but differences arise in the intermediate and dissipative ranges.
@@
 FEDONet achieves lower pointwise errors while eliminating the artificial spectral plateau that characterizes conventional fully connected trunks.
-3.2 Burgers’ Equation
-The Burgers’ equation is a well-known 1D viscous equation. It has been studied extensively for DeepONets in prior literature [53]. The equation is given by:
-$\frac{\partial u}{\partial t}(x,t)+u\frac{\partial u}{\partial x}(x,t)=\nu\frac{\partial^{2}u}{\partial x^{2}}(x,t),\>\>\>\>\>\forall(x,t)\in[0,1]\times[0,1]$ (13)
-where $x$ and $t$ denote the spatio-temporal coordinates and $\nu$ is the kinematic viscosity. We set $\nu=0.01$ along with initial and periodic boundary conditions as:
-$\begin{split}\displaystyle u(x,0)=&\>s(x),\>\>\>\>\>\>\>\>\>\>\>\forall\>x\in[0,1]\\ u(0,t)=&\>u(1,t),\>\>\>\>\>\>\>\>\forall\>t\in[0,1]\\ \frac{\partial u}{\partial x}(0,t)=&\>\frac{\partial u}{\partial x}(1,t)\>\>\>\>\>\>\forall\>t\in[0,1]\end{split}$ (14)
-where the initial condition $s(x)$ is sampled from a Gaussian Random Field, $s(x)\sim\mathcal{N}\left(0,25^{2}(-\Delta+5^{2}I)^{-4}\right)$, satisfying periodicity.
-The objective is to learn the nonlinear operator $\mathcal{G}$ mapping $s(x)$ to the full spatio-temporal solution $u(x,t)$ using both vanilla DeepONet and the proposed FEDONet.
-Both models were evaluated on $250$ unseen test samples, yielding average relative $L^{2}$ errors of $6.43\%$ for the DeepONet and $4.86\%$ for the FEDONet.
-Figure 4 presents prediction for a test sample, showing the ground truth solution alongside predictions from both models.
-While DeepONet captures the coarse solution structure, it struggles with steep gradients and exhibits visible smoothing near shock regions.
-In contrast, FEDONet produces significantly sharper and more accurate reconstructions, faithfully capturing shock fronts and fine-scale dynamics.
-Figure 4: Solving a parametric Burgers’ equation: Comparison of predicted spatio-temporal field for a random initial condition. Relative $L^{2}$ error: $4.09\%$ for DeepONet, $1.49\%$ for FEDONet.
-The energy spectra of the predicted and ground truth fields, shown in Figure 5, further emphasize this advantage.
-FEDONet accurately captures the spectral decay across a wide range of wavenumbers, aligning with the physical expectations of dissipative transport phenomena.
-DeepONet, however, underestimates energy at higher modes, reflecting its inability to resolve fine features and spectral richness.
-This demonstrates that the Fourier embeddings effectively bridge the model’s spectral bias, enabling it to better represent high-frequency dynamics essential to Burgers-type flows.
-The enhanced performance of FEDONet in both physical and spectral domains highlights its potential as a more robust and generalizable operator learner for convection-dominated systems.
-Figure 5: Energy spectrum and pointwise absolute error across space and time for a representative sample for Burgers’ Equation.
-Figure 6 shows the convergence of training losses for different batch sizes.
-FEDONet consistently converges faster and to lower loss values across all training configurations, indicating improved optimization characteristics and stability during training.
-Figure 6: Training loss convergence for the 1D Burgers’ Equation - Comparison of DeepONet and FEDONet trained with varying batch sizes.
-3.3 Lorenz-63 System
-The Lorenz-63 system is a widely studied chaotic dynamical system that serves as a benchmark for evaluating the long-term stability and predictive accuracy of surrogate models.
-It is governed by the following system of coupled nonlinear ordinary differential equations:
-$\frac{dx}{dt}$ $=\sigma(y-x),$ (15)
-$\frac{dy}{dt}$ $=x(\rho-z)-y,$
-$\frac{dz}{dt}$ $=xy-\beta z,$
-where the canonical parameters are set to $\sigma=10$, $\rho=28$, and $\beta=8/3$. Under these settings, the system exhibits chaotic dynamics on a strange attractor in three-dimensional phase space.
-The objective is to approximate the solution operator $\mathcal{G}\colon x_{0}\mapsto\mathbf{x}(t)$, where $\mathbf{x}(t)=(x(t),y(t),z(t))$ denotes the full trajectory over the time interval $t\in[0,3]$, given a scalar initial condition $x_{0}\in[10,15]$, with $y_{0}=z_{0}=12$ held fixed.
-This configuration reduces the input to a one-dimensional function-to-function mapping, while preserving the rich, sensitive behavior of the output trajectory.
-A dataset of 10,00010{,}000 trajectories is generated using the classic Runge-Kutta method (RK45), each evaluated at 1000 uniformly spaced time steps.
-The input to the DeepONet branch network is the scalar $x_{0}$, and the trunk network is queried at time coordinates $t\in[0,3]$.
-The goal is to accurately recover the full state trajectory, including lobe transitions and fine-scale oscillations characteristic of chaotic dynamics.
-Figure 7: Reconstructed Lorenz-63 attractor for a randomnly sampled initial condition for trained DeepONet and FEDONet
-Figure 7 shows the reconstructed attractor for a representative test sample.
-While DeepONet departs from the ground truth early in the rollout, FEDONet maintains geometric coherence and captures both wings of the attractor.
-This result demonstrates the stabilizing effect of spectral enrichment on trajectory learning.
-Figure 8: Phase-plane projection of the xxzz states.
-In Figure 8, we examine the xxzz projection.
-The DeepONet prediction deviates from the true phase-space trajectory and fails to capture the correct lobe transitions.
-In contrast, FEDONet successfully recovers the attractor’s phase geometry, indicating more robust learning of long-term dependencies.
-Figure 9: Pointwise prediction error in $x(t)$, $y(t)$, and $z(t)$ for a representative sample.
-Figure 9 presents the pointwise prediction errors over time.
-DeepONet accumulates noticeable errors, particularly in the $x(t)$ and $z(t)$ components, which undergo rapid transitions between lobes.
-These deviations grow sharply after $t>1.5$, illustrating compounding approximation errors.
-FEDONet exhibits substantially lower and smoother error profiles, with better robustness during nonlinear state transitions.
-Figure 10: Cumulative relative error over time.
-In Figure 10, we observe that the cumulative error of DeepONet grows exponentially, characteristic of chaotic divergence.
-In contrast, FEDONet maintains a flatter cumulative error curve, demonstrating enhanced resilience against long-term instability and better preservation of trajectory fidelity.
-Figure 11: Spectral energy density of predicted $x(t)$ trajectories compared with the ground truth.
-Finally, Figure 11 compares the spectral content of the predicted and reference trajectories.
-The vanilla DeepONet fails to capture high-frequency modes, resulting in a smoothed trajectory that diverges from the correct dynamics.
-FEDONet, equipped with randomized Fourier features, preserves the broadband energy spectrum more accurately - an essential factor in recovering the fine-scale structure and temporal richness of chaotic systems.
-These results collectively demonstrate that FEDONet offers improved phase alignment, lower pointwise and cumulative error, and superior spectral accuracy in chaotic dynamical systems.
-The introduction of Fourier embeddings into the trunk network acts as a spectral preconditioner, enabling the model to overcome inductive biases and improve fidelity under sensitivity to initial conditions.
+The robustness of the learned operators to noisy inputs is examined in Fig.
+8, which shows the variation of mean relative $L^{2}$ error with increasing levels of additive noise.
+While both models exhibit degradation in performance as noise increases, FEDONet consistently achieves lower error across all noise levels.
+Moreover, the rate of error growth is significantly smaller for FEDONet, indicating enhanced stability.
+Notably, even at higher noise levels (up to $20\%$), FEDONet maintains a relatively low error compared to DeepONet, highlighting its ability to preserve the underlying solution structure despite corrupted inputs.
+Figure 8: Robustness to input noise for the 2D Poisson equation: Mean relative $L^{2}$ error as a function of input noise level.
 3.3 Eikonal Equation
-We consider the two-dimensional Eikonal equation, which governs the propagation of wavefronts at unit speed and arises in a wide range of applications including geometric optics, level-set methods, and computational geometry.
-The equation is given by:
+We consider the 2D Eikonal equation, which governs the propagation of wavefronts at unit speed and arises in a wide range of applications including geometric optics, level-set methods, and computational geometry.
+We consider the governing equation,
 $\|\nabla s(\mathbf{x})\|_{2}$ $=1,$ (14)
@@
 where $\mathbf{x}=(x,y)\in\mathbb{R}^{2}$ denotes the spatial coordinates, $\Omega\subset\mathbb{R}^{2}$ is the computational domain, and $\partial\Omega$ is its boundary.
-The solution $s(\mathbf{x})$ represents the signed distance from a point $\mathbf{x}\in\Omega$ to the boundary $\partial\Omega$, defined as:
+The solution $s(\mathbf{x})$ represents the signed distance from a point $\mathbf{x}\in\Omega$ to the boundary $\partial\Omega$, defined as,
 […]
-with the distance function $d$ defined by:
+with the distance function $d$ defined by,
 $d(\mathbf{x},\partial\Omega):=\inf_{\mathbf{y}\in\partial\Omega}\|\mathbf{x}-\mathbf{y}\|_{2}.$ (16)
@@
 As a testbed, we construct a dataset of NACA 4-digit airfoil shapes, widely used in aerodynamics, and compute their SDFs within a structured computational domain.
-A NACA 4-digit airfoil is parameterized by its maximum camber $m$, the location of maximum camber $p$, and the maximum thickness $t$. The mean camber line is given by:
+A NACA 4-digit airfoil is parameterized by its maximum camber $m$, the location of maximum camber $p$, and the maximum thickness $t$. The mean camber line is given by,
 $y_{c}(x)=\begin{cases}\frac{m}{p^{2}}(2px-x^{2}),&0\leq x<p,\\ \frac{m}{(1-p)^{2}}((1-2p)+2px-x^{2}),&p\leq x\leq 1,\end{cases}$ (17)
-and its derivative is:
+and its derivative is,
 $\frac{dy_{c}}{dx}=\begin{cases}\frac{2m}{p^{2}}(p-x),&0\leq x<p,\\ \frac{2m}{(1-p)^{2}}(p-x),&p\leq x\leq 1.\end{cases}$ (18)
-The thickness distribution $y_{t}(x)$ is defined as:
+The thickness distribution $y_{t}(x)$ is defined as,
 $y_{t}(x)=5t\left(0.2969\sqrt{x}-0.1260x-0.3516x^{2}+0.2843x^{3}-0.1036x^{4}\right).$ (19)
-The upper and lower surface coordinates of the airfoil are computed as:
+The upper and lower surface coordinates of the airfoil are computed as,
 $x_{u}=x-y_{t}\sin(\theta),\quad$ $y_{u}=y_{c}+y_{t}\cos(\theta),$ (20)
@@
 Each airfoil is embedded in a $256\times 256$ pixel computational grid and scaled to fit within a central subdomain.
-A binary mask is created to define the geometry, and the signed distance function is computed using the Euclidean Distance Transform (EDT) as:
+A binary mask is created to define the geometry, and the signed distance function is computed using the Euclidean Distance Transform (EDT) as,
 $s(x,y)=d_{\text{out}}(x,y)-d_{\text{in}}(x,y),$ (22)
 where $d_{\text{out}}$ and $d_{\text{in}}$ represent the unsigned distances outside and inside the airfoil boundary, respectively.
-To ensure numerical stability, the resulting SDF is normalized to the range [1,1][-1,1]:
+To ensure numerical stability, the resulting SDF is normalized to the range [1,1][-1,1],
 $s_{\text{norm}}(x,y)=\frac{s(x,y)}{\max|s(x,y)|}.$ (23)
@@
 This dataset captures a wide range of geometrical variations and curvature profiles, providing a rigorous benchmark for learning PDE solution operators from geometric input.
-On this benchmark, the baseline DeepONet attains an average relative $L^{2}$ error of 2.15%, while the proposed FEDONet achieves an improved 1.12% error, demonstrating substantially higher reconstruction accuracy.
-Figure 12: Solving a parametric Eikonal equation (airfoils): Comparison of predicted signed distance functions (SDFs) for a representative airfoil geometry.
-Relative $L^{2}$ error: $2.728\times 10^{-2}$ for DeepONet, $0.912\times 10^{-2}$ for FEDONet.
+On this benchmark, the baseline DeepONet attains an average relative $L^{2}$ error of 2.15%, while the proposed FEDONet achieves an improved 1.12% error under noise-free input conditions, demonstrating substantially higher reconstruction accuracy.
+Figure 9: Solving a parametric Eikonal equation (airfoils): Comparison of predicted signed distance functions (SDFs) for a representative noise-free airfoil geometry.
+The corresponding relative $L^{2}$ errors are $2.34\%$ for DeepONet and $0.54\%$ for FEDONet.
 As shown in Figure 9, FEDONet achieves more accurate recovery of the level-set structure.
@@
 FEDONet’s use of Fourier embeddings in the trunk network enhances its ability to resolve these features by expanding the function space with oscillatory basis functions aligned with the data’s spectral demands.
+This advantage is further reflected in the structure of the error fields, where the DeepONet solution exhibits anisotropic, ray-like artifacts emanating from regions of high curvature, indicative of an uneven representation of spatial frequencies and a directionally biased approximation of the underlying operator.
+Such behavior is consistent with the spectral bias of standard neural networks, which tend to under-resolve high-frequency components.
+In contrast, FEDONet produces a more isotropic and spatially coherent error distribution with reduced magnitude, demonstrating improved consistency across scales and a more faithful reconstruction of the solution, particularly in regions dominated by sharp gradients and geometric complexity.
+The effect of training dataset size on model performance is shown in Fig.
+10.
+Both models benefit from increased training data; however, FEDONet consistently achieves lower mean relative $L^{2}$ error across all sample sizes.
+The improvement is particularly significant in the low-data regime, which indicates that the Fourier-embedded architecture is more sample efficient and better suited for learning the nonlinear mapping associated with the Eikonal equation, which involves sharp geometric features and non-smooth solution manifolds.
+Figure 10: Sample efficiency for the Eikonal equation: Mean relative $L^{2}$ error as a function of the number of training samples.
+FEDONet consistently achieves lower error across all regimes, with a pronounced advantage in the low-data setting.
+Figure 11 presents qualitative comparisons of model predictions under increasing levels of additive noise in the input mask.
+As the noise level increases, the DeepONet predictions exhibit pronounced degradation, with large, structured error patterns emerging near regions of high curvature and along characteristic directions.
+These artifacts manifest as anisotropic streaks and localized error amplification, indicating instability in the learned operator under perturbations.
+In contrast, FEDONet maintains stable and visually consistent reconstructions across all noise levels.
+The corresponding error fields remain relatively diffuse and isotropic, with significantly reduced magnitude even at higher noise levels.
+Figure 11: Robustness to input noise for the Eikonal equation: Qualitative comparison of predicted fields and corresponding error distributions under increasing noise levels.
 These results confirm that FEDONet offers improved resolution of high-gradient and discontinuous regions in PDE fields, even in geometry-conditioned tasks such as the Eikonal equation.
-This highlights its potential as a general-purpose surrogate for PDEs defined on complex, spatially varying domains.
-3.5 Allen–Cahn Equation
-We further evaluate the performance of the proposed FEDONet in modeling the one-dimensional Allen–Cahn nonlinear reaction-diffusion system widely used to describe phase separation processes in multi-component systems.
-The governing partial differential equation is given by:
+This demonstrates that the Fourier embeddings provide enhanced robustness by enabling a more uniform representation of spatial frequencies, thereby preserving the underlying geometric structure of the solution.
+3.4 Allen-Cahn Equation
+We further evaluate the performance of FEDONet in modeling the 1D Allen-Cahn nonlinear reaction-diffusion system, widely used to describe phase separation processes in multi-component systems.
+The governing partial differential equation is given by,
 $\frac{\partial u}{\partial t}=\epsilon\frac{\partial^{2}u}{\partial x^{2}}-5u^{3}+5u,\quad x\in[-1,1],\quad t\in[0,1],$ (24)
@@
 The spatial and temporal grids are discretized using $\Delta x=0.01$ and $\Delta t=0.005$, yielding solution arrays of dimension $200\times 200$.
-Initial conditions are defined as:
+Initial conditions are defined as,
 $s(x)=\sum_{k=1}^{3}\left[a_{k}x^{2k}\cos(k\pi x)+b_{k}x^{2k}\sin(k\pi x)\right],$ (25)
@@
 The goal is to approximate the solution operator $\mathcal{G}\colon s(x)\mapsto u(x,t)$, comparing the performance of vanilla DeepONet and FEDONet.
-Figure 13: Solving a parametric Allen Cahn equation: Comparison of predicted spatio-temporal field for a random initial condition. Relative $L^{2}$ error: $4.85\%$ for DeepONet, $2.27\%$ for FEDONet.
-Both models were evaluated on $128$ randomly selected unseen test samples, yielding average relative $L^{2}$ errors of $10.61\%$ for the DeepONet and $5.50\%$ for the FEDONet.
-Figure 13 shows qualitative predictions for a representative test case.
-FEDONet sharply resolves the interface zones and phase transitions, while DeepONet produces smoother approximations with notable blurring near discontinuities.
-The pointwise absolute error plot in Figure 14 emphasizes the localized improvements of FEDONet in capturing steep interfaces.
-DeepONet accumulates substantial errors in regions with high spatial gradients, whereas FEDONet maintains predictive accuracy throughout the domain, suggesting better handling of dissipative nonlinearities.
-Figure 14: Energy spectrum and pointwise absolute error across space and time for a representative sample for Allen-Cahn Equation.
-Figure 14 compares the spatial energy spectra of predicted and true solutions.
-FEDONet accurately tracks the ground truth spectrum across all wavenumbers, preserving the high-frequency tail associated with sharp phase boundaries.
-In contrast, DeepONet suffers from spectral attenuation at $k>20$, revealing an inability to fully capture dissipative and oscillatory dynamics inherent in the Allen-Cahn system.
-3.6 Kuramoto–Sivashinsky Equation
+Figure 12: Solving a parametric Allen-Cahn equation: Comparison of predicted spatio-temporal solution fields for a test input with additive noise.
+The corresponding relative $L^{2}$ errors are $7.64\%$ for DeepONet and $3.16\%$ for FEDONet.
+Both models were evaluated on $250$ randomly selected, unseen, noise-free test samples, yielding average relative $L^{2}$ errors of $10.98\%$ for DeepONet and $5.85\%$ for FEDONet.
+Figure 12 presents qualitative predictions for a representative test case with $10\%$ additive noise in the input.
+While DeepONet captures the overall structure of the solution, it exhibits noticeable smoothing and loss of sharpness near interface regions.
+In contrast, FEDONet produces significantly sharper reconstructions, accurately resolving interface zones and phase transitions with minimal blurring.
+Figure 13: Spectral and Error analysis for Allen-Cahn equation: Comparison of the energy spectrum and pointwise absolute error fields for a representative noise-free test sample.
+The pointwise absolute error profiles for a representative noise-free test case, shown in Figure 13, highlight the localized accuracy of FEDONet in resolving steep interface regions.
+DeepONet exhibits significant error accumulation in regions with sharp spatial gradients, particularly near evolving interfaces, reflecting its tendency to under-resolve high-frequency components.
+In contrast, FEDONet maintains consistently low error across the domain, with only minor localized deviations, indicating a superior ability to capture sharp transitions and preserve interface structure.
+This behavior is further corroborated by the spectral analysis in Figure 13, where the energy spectra of the predicted and true solutions are compared.
+FEDONet closely tracks the ground truth spectrum, whereas DeepONet exhibits noticeable spectral attenuation for $k>20$.
+3.5 Kuramoto-Sivashinsky Equation
 We evaluate the performance of FEDONet on the one-dimensional Kuramoto-Sivashinsky (KS) equation, a canonical model that exhibits spatiotemporal chaos due to the combined effects of nonlinear advection, destabilizing diffusion, and stabilizing fourth-order hyperviscosity.
-The governing equation is given by:
+The governing equation is given by,
 $\begin{split}&\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}+\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{4}u}{\partial x^{4}}=0,\quad x\in[0,L],\;t\in[0,T],\\ &u(x,0)=u_{0}(x),\quad x\in[0,L],\end{split}$ (26)
-where $u(x,t)$ is the scalar field of interest, and $L=24$, $T=50$ denote the spatial and temporal domains respectively. The initial condition $u_{0}(x)$ is synthesized as a randomized Fourier series:
+where $u(x,t)$ is the scalar field of interest, and $L=24$, $T=50$ denote the spatial and temporal domains, respectively.
+The initial condition $u_{0}(x)$ is synthesized as a randomized Fourier series,
 $u_{0}(x)=\sum_{n=1}^{4}C_{n}\sin\left(\frac{nx}{L}\right),\quad C_{n}\sim\mathcal{N}(0,1),$ (27)
@@
 We generate 10,000 such initial conditions and simulate the corresponding solution fields using a pseudo-spectral solver.
-The objective is to learn the operator $\mathcal{G}\colon u_{0}(x)\mapsto u(x,t)$ using both the standard DeepONet and the proposed FEDONet models.
+The resulting spatial and temporal grids are subsequently normalized to the unit domain $0,1]\times[0,1]$ for operator learning.
+The objective is to learn the operator $\mathcal{G}\colon u_{0}(x)\mapsto u(x,t)$ using both the standard DeepONet and the FEDONet.
 This problem serves as a stringent benchmark due to the sensitivity of chaotic systems to initial conditions and their broadband frequency content.
-Figure 15: Solving a parametric KS equation: Median performing test-sample. Relative $L^{2}$ error: FEDONet = 16.36%, DeepONet = 89.63%.
+Figure 14: Solving a parametric KS equation: Median performing test-sample. The corresponding relative $L^{2}$ errors are 89.63% for DeepONet and 16.36% for FEDONet.
 Figure 14 shows the reconstructed spatiotemporal field for a median test case.
@@
 In stark contrast, the vanilla DeepONet produces incoherent and smoothed reconstructions, with pronounced phase lag and energy collapse.
-This degradation is reflected in a high relative error of $89.63\%$ compared to just $16.36\%$ for FEDONet.
+This degradation is reflected in a high average relative error of $89.63\%$ compared to just $16.36\%$ for FEDONet for 512 unseen noise-free test samples.
+Figure 15: Solving a parametric KS equation: Best performing test sample.
+FEDONet retains the amplitude and frequency content of the chaotic trajectory, yielding a low relative $L^{2}$ error of $10.77\%$.
 In Figure 15, we highlight the best-performing test case.
@@
 Notably, the trajectory remains consistent with the ground truth over long integration times, demonstrating that the spectral trunk can stabilize predictions and suppress drift.
-Such performance underscores the model’s ability to retain energy in high-frequency modes essential for modeling non-periodic fine-grained behavior that typifies chaotic fields.
-Figure 16: Solving a parametric KS equation: Best performing test sample.
-FEDONet retains the amplitude and frequency content of the chaotic trajectory, yielding a low relative $L^{2}$ error of $10.77\%$.
-Figure 17: Solving a parametric KS equation: Worst performing test sample.
-Even in failure cases, FEDONet remains qualitatively reasonable.
-Relative $L^{2}$ error: FEDONet = $58.02\%$, versus near-complete divergence for DeepONet.
-Figure 17 displays the worst-case test result.
-While the FEDONet prediction exhibits growing deviations from the reference field, it still retains discernible wave structures and phase coherence, with an error of $58.02\%$.
-In comparison, the vanilla DeepONet fails entirely, producing numerical artifacts, amplitude distortion, and full trajectory collapse.
-This robustness under extreme conditions emphasizes the value of frequency-enhanced representations in learning dynamical system operators.
-Figure 18 plots the training loss curves for both models.
-FEDONet converges faster and exhibits more stable optimization behavior, attributable to the improved conditioning of the input space via randomized Fourier features.
-This aligns with observations from kernel learning and feature engineering literature, where frequency-based transformations are known to facilitate smoother descent landscapes and gradient propagation.
-Figure 18: KS Training convergence comparison. FEDONet achieves faster and more stable optimization.
-Figure 19: Distribution of relative $L^{2}$ errors across 512 unseen test samples for KS Equation.
-FEDONet maintains lower error in both mean and variance, demonstrating improved robustness and generalization in chaotic regimes.
-Finally, Figure 19 summarizes performance statistics across 512 randomly selected test cases.
-FEDONet consistently achieves lower relative $L^{2}$ errors with reduced variance, indicating better generalization and resilience to initial condition perturbations.
-In totality, the Kuramoto-Sivashinsky benchmark clearly illustrates the limitations of standard DeepONet trunks in handling chaotic, broadband dynamics.
-The introduction of Fourier embeddings into the trunk network provides an effective inductive bias that promotes spectral alignment, temporal coherence, and long-horizon accuracy.
-FEDONet emerges as a compelling model for operator learning in complex regimes, particularly those characterized by sensitive dependence on initial conditions and multiscale feature interactions.
+Figure 16: Robustness to input noise for the KS equation: Qualitative comparison of predicted spatiotemporal fields under increasing noise levels in the initial condition.
+The quantitative impact of input noise is summarized in Fig.
+16.
+DeepONet exhibits consistently high error even in the noise-free setting and shows only marginal variation with increasing noise, indicating a fundamental inability to capture the underlying operator.
+In contrast, FEDONet achieves substantially lower error across all noise levels, although a gradual increase is observed as noise increases.
+Notably, even at high noise levels (up to $30\%$), FEDONet significantly outperforms DeepONet, demonstrating superior robustness and stability in chaotic regimes.
 4 Summary
-In this work, we propose the Fourier-Embedded DeepONet (FEDONet) as a principled extension to the classical DeepONet architecture for learning nonlinear operators governed by partial differential equations.
-By injecting randomized Fourier features into the trunk input, FEDONet expands the hypothesis space of DeepONets, offering enhanced spectral representation, improved numerical conditioning, and faster convergence during training all without increasing runtime complexity.
-FEDONet consistently outperforms the vanilla DeepONet.
-Notably, it delivers dramatic reductions in relative $L^{2}$ error, most prominently in multiscale and high-frequency scenarios such as the Kuramoto-Sivashinsky equation (76% $\rightarrow$ 21%).
-Table 1: Relative $L^{2}$ error (%) for each benchmark.
-Dataset DeepONet FEDONet
-2-D Poisson (elliptic) 1.41 1.09
-1-D Burgers (conv./diff.) 6.43 4.86
-Lorenz-63 (chaotic ODE) 2.03 0.46
-Allen-Cahn (phase–field) 10.61 5.50
-Kuramoto-Sivashinsky (spatio-chaos) 75.11 20.62
-Eikonal (Hamilton–Jacobi) 2.15 1.12
-Beyond pointwise accuracy, FEDONet preserves physical consistency across modalities.
-Energy spectrum analyses reveal that FEDONet better captures inertial and dissipation range behavior, avoiding the artificial spectral pile-up observed in DeepONets.
-In geometry-sensitive tasks like the Eikonal equation, it reconstructs sharp features and discontinuities with higher fidelity.
-Taken together, these results establish FEDONet as a robust, scalable, and architecture-agnostic enhancement for neural operator learning.
-The proposed spectral trunk serves as a drop-in replacement that aligns the inductive biases of DeepONets with the spectral demands of complex PDE systems, providing a promising path forward for efficient surrogate modeling in scientific machine learning.
-5 Future Work
-The Fourier-Embedded DeepONet (FEDONet) architecture presents a compelling foundation for enhancing operator learning, yet several promising extensions remain unexplored.
-A natural next step is to generalize beyond sinusoidal embeddings by investigating alternative basis functions in the trunk network.
-Wavelets, Chebyshev polynomials, or orthogonal polynomials tailored to specific PDE families could offer improved localization or boundary adaptivity, especially for systems exhibiting sharp interfaces or multi-scale dynamics.
-Additionally, hybridizing spectral embeddings with structured architectures such as graph-based trunks or encoder-decoder networks may further boost performance on irregular domains or hierarchical solution spaces.
-Another avenue involves making the Fourier embedding parameters learnable.
-While the current design uses fixed random frequency matrices, enabling adaptive tuning of these frequencies could allow the model to better align with the intrinsic spectral content of the data.
-This modification could yield a more expressive and data-efficient hypothesis space while preserving the training stability benefits of approximate whitening.
-Finally, incorporating uncertainty quantification and physics-informed constraints into the FEDONet framework via Bayesian formulations or PINO style residual losses may further enhance robustness, interpretability, and generalization in scientific computing tasks involving sparse or noisy data.
-Declaration of competing interest
+In this work, we introduced the Fourier-Embedded DeepONet (FEDONet), a principled extension of the classical DeepONet architecture for learning nonlinear operators governed by partial differential equations.
+By applying randomized Fourier feature embeddings to the trunk coordinates, FEDONet performs a spectral lifting of the input space that can be interpreted as a form of spectral preconditioning, improving the conditioning of the learning problem, expanding the effective hypothesis space, and mitigating the spectral bias of standard multilayer perceptrons.
+This enables efficient representation of multiscale and oscillatory solution structures while preserving the original branch-trunk formulation and incurring negligible computational overhead.
+Extensive evaluations across a diverse suite of canonical PDE benchmarks demonstrate that FEDONet consistently outperforms the vanilla DeepONet across elliptic, parabolic, hyperbolic, and chaotic regimes, with particularly significant gains in systems exhibiting broadband frequency content such as the Kuramoto-Sivashinsky equation.
+A comprehensive summary of the results is provided in Table 1.
+Beyond pointwise accuracy, FEDONet exhibits superior spectral fidelity, accurately capturing inertial and dissipative ranges without artificial high-frequency energy accumulation, and demonstrates enhanced robustness to input noise as well as improved sample efficiency in low-data regimes.
+Unlike global spectral operator methods such as Fourier Neural Operators, FEDONet retains the locality and modularity of the DeepONet framework while introducing a principled spectral inductive bias, establishing it as a simple, scalable, and architecture-agnostic enhancement for high-fidelity operator learning in scientific machine learning.
+Table 1: Mean relative $L^{2}$ error (%) of FEDONet and DeepONet.
+Dataset DeepONet FEDONet Improvement
+1-D Burgers (convection-diffusion) 6.43 4.86 24.4%
+2-D Poisson (elliptic) 1.41 1.09 22.7%
+Eikonal (Hamilton-Jacobi) 2.15 1.12 47.9%
+Allen-Cahn (phase-field) 10.98 5.85 46.72%
+Kuramoto-Sivashinsky (spatio-temporal chaos) 89.63 16.36 81.75%
+5 Conclusion and Future Work
+While FEDONet demonstrates strong performance across a range of operator learning tasks, several directions remain for further development, particularly toward real-world applicability.
+A natural extension is to move beyond fixed sinusoidal embeddings by exploring alternative spectral or localized basis functions, such as wavelets or orthogonal polynomials, to improve spatial localization and boundary adaptivity.
+Integrating the spectral embedding strategy with structured architectures, including graph-based trunks or multiscale encoder-decoder frameworks, also offers a pathway toward handling complex geometries and irregular domains.
+Another promising direction is the introduction of adaptive or learnable spectral embeddings, which could better align with the intrinsic frequency content of the underlying operator and improve data efficiency.
+In addition, incorporating uncertainty quantification and physics-based constraints will be essential for robust deployment in practical settings.
+Although this study focuses on canonical PDE benchmarks, FEDONet is motivated by challenges in high-dimensional systems such as turbulent flows and geophysical dynamics, where multiscale structure, noise sensitivity, and spectral complexity are dominant.
+The Fourier-embedded trunk provides a lightweight mechanism for introducing spectral inductive bias, making the framework well-suited for applications in surrogate modeling and real-time prediction.
+Key challenges remain, including scalability to high-resolution domains, handling complex geometries, and integration with data assimilation or hybrid physics-informed frameworks.
+Addressing these will be critical for extending operator learning methods from controlled benchmarks to realistic scientific applications.
+Declaration of Competing Interests
 The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@
 Data Availability
-Data supporting the findings of this study are available from the corresponding author upon reasonable request.
+The datasets and code used to generate the results presented in this study are publicly available and can be accessed through the following GitHub repository: https://github.com/as26101999/Fourier-Embedded-DeepONets
 Declaration of Generative AI use
@@
 Let $\phi(\zeta)=\sqrt{2}\,[\sin(2\pi Z\zeta),\cos(2\pi Z\zeta)]\in\mathbb{R}^{2M}$ be the Fourier feature embedding of an input $\zeta\in\mathbb{R}^{d}$, where each row of the matrix $Z\in\mathbb{R}^{M\times d}$ is sampled i.i.d.
-from a Gaussian distribution:
+from a Gaussian distribution,
 $Z_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (28)
@@
 $\mathbb{E}_{\zeta}[\phi(\zeta)\phi(\zeta)^{\top}]\approx I_{2M}.$ (29)
-To see this, consider a single frequency row vector $Z_{i}\in\mathbb{R}^{1\times d}$ drawn from $Z$. The corresponding $2$-dimensional block of the embedding is:
+To see this, consider a single frequency row vector $Z_{i}\in\mathbb{R}^{1\times d}$ drawn from $Z$. The corresponding $2$-dimensional block of the embedding is,
 $\phi_{i}(\zeta)=\sqrt{2}\begin{bmatrix}\sin(2\pi Z_{i}\cdot\zeta)\\ \cos(2\pi Z_{i}\cdot\zeta)\end{bmatrix}$. (30)
 We now compute the second-order statistics of this embedding block under the assumption that $Z_{i}\cdot\zeta$ is uniformly distributed over $[0,1]$ (a reasonable approximation due to the randomness in $Z_{i}$ and the uniformity of $\zeta$).
-Using standard trigonometric integrals, we obtain:
+Using standard trigonometric integrals, we obtain,
 $\mathbb{E}_{\zeta}[\sin^{2}(2\pi Z_{i}\cdot\zeta)]$ $\approx\frac{1}{2},$ (31)
@@
 $\mathbb{E}_{\zeta}[\sin(2\pi Z_{i}\cdot\zeta)\cos(2\pi Z_{i}\cdot\zeta)]$ $\approx 0.$ (33)
-Therefore, the unscaled covariance matrix for this block is approximately:
+Therefore, the unscaled covariance matrix for this block is approximately,
 $\mathbb{E}_{\zeta}\left[\begin{bmatrix}\sin(2\pi Z_{i}\cdot\zeta)\\ \cos(2\pi Z_{i}\cdot\zeta)\end{bmatrix}\begin{bmatrix}\sin(2\pi Z_{i}\cdot\zeta)&\cos(2\pi Z_{i}\cdot\zeta)\end{bmatrix}\right]\approx\frac{1}{2}I_{2}.$ (34)
-Scaling by the $\sqrt{2}$ factor yields:
+Scaling by the $\sqrt{2}$ factor yields,
 $\mathbb{E}_{\zeta}[\phi_{i}(\zeta)\phi_{i}(\zeta)^{\top}]\approx I_{2}.$ (35)
-Since the rows of $Z$ are sampled independently and the embeddings for each $Z_{i}$ are uncorrelated, the full embedding vector $\phi(\zeta)\in\mathbb{R}^{2M}$ satisfies:
+Since the rows of $Z$ are sampled independently and the embeddings for each $Z_{i}$ are uncorrelated, the full embedding vector $\phi(\zeta)\in\mathbb{R}^{2M}$ satisfies,
 $\mathbb{E}_{\zeta}[\phi(\zeta)\phi(\zeta)^{\top}]\approx I_{2M}.$ (36)
@@
 As such, Fourier embeddings not only enhance expressivity but also promote well-conditioned training dynamics.
-Appendix B Theoretical Justification for Fourier Embedding Superset Property
-In this appendix, we formally justify the claim stated in Section 2.1 that the hypothesis space induced by Fourier embeddings strictly contains the vanilla hypothesis space of coordinate-input networks.
-We present this as a lemma and provide constructive proof using classical function approximation theory.
-Let $\mathcal{H}_{\text{vanilla}}$ be the hypothesis space of functions representable by a feedforward neural network with input $\zeta\in\mathbb{R}^{d}$, and let $\mathcal{H}_{\phi}$ be the hypothesis space of functions representable by a similar network where the input is first mapped via the Fourier feature embedding:
-$\phi(\zeta)=\left[\sin(2\pi Z\zeta),\cos(2\pi Z\zeta)\right],\quad Z\in\mathbb{R}^{M\times d},\;Z_{ij}\sim\mathcal{N}(0,\sigma^{2}).$ (39)
-Then, under mild regularity assumptions (e.g., compact input domain $K\subset\mathbb{R}^{d}$ and continuous activation functions), we have:
-$\mathcal{H}_{\text{vanilla}}\subsetneq\mathcal{H}_{\phi}.$ (40)
-We divide the proof into two parts.
-Part I: Inclusion. Let $f\in\mathcal{H}_{\text{vanilla}}$ be a function representable by a neural network of the form:
-$f(\zeta)=\sum_{j=1}^{N}a_{j}\,\sigma(w_{j}^{\top}\zeta+b_{j}).$ (41)
-Since each $\zeta_{i}$ (coordinate function) can be approximated by a Fourier series on a compact domain $K$, and the Fourier embedding $\phi(\zeta)$ contains these sinusoidal components, it follows that a function of $\zeta$ can be emulated by a function of $\phi(\zeta)$.
-In particular, for every $f\in\mathcal{H}_{\text{vanilla}}$ and $\varepsilon>0$, there exists $\tilde{f}\in\mathcal{H}_{\phi}$ such that:
-$\sup_{\zeta\in K}|f(\zeta)-\tilde{f}(\zeta)|<\varepsilon.$ (42)
-Therefore, $\mathcal{H}_{\text{vanilla}}\subseteq\mathcal{H}_{\phi}$.
-Part II: Strict Containment.
-Consider the function $f(\zeta)=\sin(100\pi\zeta)$ over $\zeta\in[0,1]$.
-Due to the spectral bias of neural networks [51], this high-frequency function is difficult to approximate using shallow networks on raw input.
-However, if the frequency matrix $Z$ used in the embedding $\phi(\zeta)$ includes rows close to $100$, then:
-$\sin(2\pi Z_{j}\zeta)\approx\sin(100\pi\zeta),$
-making $f$ easily representable in $\mathcal{H}_{\phi}$.
-Since $f(\zeta)$ cannot be efficiently approximated in $\mathcal{H}_{\text{vanilla}}$ (requiring significantly more parameters or depth), but is trivially accessible in $\mathcal{H}_{\phi}$, we conclude that the inclusion is strict:
-$\mathcal{H}_{\text{vanilla}}\subsetneq\mathcal{H}_{\phi}.$ (43)
-This result formalizes the intuition that Fourier embeddings expand the expressivity of the network and can be interpreted as a form of spectral preconditioning for operator learning models such as DeepONets.
-Appendix C Spectral Interpretation of Fourier-Embedded DeepONets
-The use of Fourier embeddings in the trunk network induces a natural spectral decomposition in the DeepONet output.
-Specifically, the composition of the trunk and branch networks can be interpreted as a data-driven spectral projection:
-$\mathcal{G}(u)(\zeta)\approx\sum_{k=1}^{p}\langle u,\varphi_{k}\rangle\,\psi_{k}(\zeta),$ (44)
-where $\{\varphi_{k}\}_{k=1}^{p}$ are data-dependent basis functions learned from the branch network, and $\{\psi_{k}(\zeta)\}_{k=1}^{p}$ are basis functions defined by the trunk network, shaped by the sinusoidal feature embedding $\phi(\zeta)$.
-When the trunk network is fed sinusoidal embeddings of the form
-$\phi(\zeta)=[\cos(2\pi\omega_{1}^{\top}\zeta),\dots,\cos(2\pi\omega_{M}^{\top}\zeta),\sin(2\pi\omega_{1}^{\top}\zeta),\dots,\sin(2\pi\omega_{M}^{\top}\zeta)],$
-the basis functions $\psi_{k}(\zeta)$ are learned combinations of trigonometric modes.
-This allows the model to represent oscillatory and localized patterns that are characteristic of multiscale physical systems.
-The representation in Equation (44) mirrors classical Galerkin methods, where the solution is projected onto a known orthonormal basis:
-$u(x,t)\approx\sum_{n=1}^{N}a_{n}(t)\,\phi_{n}(x),$ (45)
-with the coefficients $a_{n}(t)$ governed by PDE dynamics.
-In DeepONet, both the basis functions $\psi_{k}$ and the projection coefficients $\langle u,\varphi_{k}\rangle$ are learned directly from data, enabling generalization beyond handcrafted bases.
-Unlike classical global bases such as Fourier or Chebyshev functions, the neural trunk network can generate localized, adaptive basis functions.
-This enhances the model’s ability to approximate functions with sharp transitions, discontinuities, or compactly supported features, which are common in real-world PDE systems.
-We may formalize the DeepONet architecture as a composition of a projection and synthesis operator:
-$\Pi_{\varphi}(u):=[\langle u,\varphi_{1}\rangle,\dots,\langle u,\varphi_{p}\rangle]^{\top},\quad\mathcal{S}_{\psi}(\zeta):=[\psi_{1}(\zeta),\dots,\psi_{p}(\zeta)]^{\top},$
-yielding the output:
-$\mathcal{G}(u)(\zeta)\approx\mathcal{S}_{\psi}(\zeta)^{\top}\Pi_{\varphi}(u),$
-which resembles a spectral synthesis operation of the form:
-$f(x)\approx\sum_{k}\hat{f}_{k}\phi_{k}(x).$
-This interpretation leads to several insights:
-  1. 1.
-    Fourier-Embedded DeepONets naturally align with the dominant spectral modes of the target operator, facilitating more efficient approximation of high-frequency content.
-  2. 2.
-    The learned basis can adapt to complex patterns in the data, unlike fixed basis expansions.
-  3. 3.
-    The operator learning framework parallels well-established numerical discretization techniques, such as Petrov-Galerkin and spectral methods, but operates in a fully data-driven regime.
-In summary, Fourier Embeddings serve as a principled mechanism for spectral lifting, enabling DeepONet to perform a learned Galerkin-style decomposition of operators.
-This perspective bridges classical PDE approximation theory with modern deep learning, offering both interpretability and flexibility in learning complex function-to-function mappings.
+Appendix B Additional Results
+B.1 Burgers’ Equation
+Figure 17: Robustness to input noise for Burgers’ equation: Mean relative $L^{2}$ error as a function of input noise level.
+FEDONet consistently outperforms DeepONet and exhibits slower error degradation.
+B.2 2D Poisson Equation
+Figure 18: Sample efficiency for the 2D Poisson equation: Mean relative $L^{2}$ error as a function of the number of training samples.
+B.3 Eikonal Equation
+Figure 19: Robustness to input noise for the Eikonal equation: Mean relative $L^{2}$ error as a function of input noise level.
+FEDONet shows consistently lower error and slower degradation compared to DeepONet, indicating improved robustness.
+B.4 Allen-Cahn Equation
+Figure 20: Sample efficiency for the Allen-Cahn equation: Mean relative $L^{2}$ error as a function of the number of training samples.
+Figure 21: Robustness to input noise for Allen-Cahn equation: Mean relative $L^{2}$ error as a function of input noise level.
+B.5 Kuramoto-Sivashinsky Equation
+Figure 22: Solving a parametric KS equation: Worst performing test sample.
+Even in failure cases, FEDONet remains qualitatively reasonable.
+Relative $L^{2}$ error: FEDONet = $58.02\%$, versus near-complete divergence for DeepONet.
+Figure 23: Robustness to input noise for KS equation: Mean relative $L^{2}$ error as a function of input noise level.